### v2

# Hallucination Detection in LLMs with Topological Divergence on Attention Graphs

Abstract
Hallucination, i.e., generating factually incorrect content, remains a critical challenge for large language models (LLMs). We introduce TOHA [^1] The code of the proposed method and the considered baselines is available at https://anonymous.4open.science/r/tda4hallu-BED5, a TOpology-based HAllucination detector in the RAG setting, which leverages a topological divergence metric to quantify the structural properties of graphs induced by attention matrices. Examining the topological divergence between prompt and response subgraphs reveals consistent patterns: higher divergence values in specific attention heads correlate with hallucinated outputs, independent of the dataset. Extensive experiments — including evaluation on question answering and summarization tasks — show that our approach achieves state-of-the-art or competitive results on several benchmarks while requiring minimal annotated data and computational resources. Our findings suggest that analyzing the topological structure of attention matrices can serve as an efficient and robust indicator of factual reliability in LLMs.
[…]
Hallucination Detection in LLMs with Topological Divergence on Attention Graphs
Alexandra Bazarova1, Aleksandr Yugay1, Andrey Shulga1, Alina Ermilova1, Andrei Volodichev1, Konstantin Polev2, Julia Belikova2, Rauf Parchiev 2, Dmitry Simakov2, Maxim Savchenko2, Andrey Savchenko2, Serguei Barannikov1, 3, Alexey Zaytsev1, 1Skolkovo Institute of Science and Technology, 2Sber AI Lab, 3CNRS, Universite Paris Cite, Correspondence: a.bazarova@skoltech.ru
**footnotetext: Equal contribution.
1 Introduction
Large language models (LLMs) have progressed significantly in recent years, finding applications in various fields Chkirbene et al. (2024). However, these models are prone to generate so-called hallucinations, i.e., content that is factually or contextually incorrect Huang et al. (2023). Detecting hallucinations is crucial for safely deploying LLMs in sensitive fields since erroneous outputs may seriously harm user trust. An effective detector would therefore expand the scope of LLM applications while mitigating risks Gao et al. (2024).
Multiple methods address this problem Huang et al. (2023); Sahoo et al. (2024), though many face significant practical constraints. A common limitation is the reliance on large annotated datasets Sky et al. (2024); Azaria and Mitchell (2023); Chuang et al. (2024), which are rarely available publicly Zhang et al. (2023) and require extensive annotation effort for each new model released. Another popular approach depends on generating multiple samples for scoring Manakul et al. (2024); Chen et al. (2024); Farquhar et al. (2024), which increases computational costs substantially.
Figure 1: Difference between average topological divergence values for hallucinated and grounded samples per attention head/layer, evaluated on MS MARCO and CoQA datasets. A lighter color corresponds to a greater difference. Green frames highlight the heads that separate samples best. The same attention heads assign greater divergence values to the hallucinated samples in both datasets. Model: Mistral-7B-Instruct-v0.1.
We address these challenges by introducing TOHA (TOpology-based HAllucination detector), a training-free method for the retrieval-augmented generation (RAG) setting Gao et al. (2023). Following the prior work Du et al. (2024), TOHA requires minimal annotated data (just $50$ annotated samples suffice for reliable detection, see Figure 4) while avoiding the computational overhead of multiple generations, making it both data- and compute-efficient.
The core insight behind TOHA is that hallucinated answers tend to have weaker connections to the context than grounded ones in the RAG setting. TOHA formalizes this by analyzing attention graphs — complete graphs induced by LLM attention maps, a representation previously used in other NLP tasks Kushnareva et al. (2021); Tulchinskii et al. (2023). Unlike prior works that rely on simplistic classifiers over basic graph properties Proskurina et al. (2023a); Cherniavskii et al. (2022), TOHA advances the approach by computing the topological dissimilarity between the attention subgraphs of the model’s response and the given context. This dissimilarity, measured via our adaptation of Manifold Topology Divergence  Barannikov et al. (2021) for the graph setting, quantifies the “strength” of context-response ties: higher values indicate weaker links and thus likely hallucinations. We prove key stability properties for this metric, ensuring its reliability as a hallucination score.
Figure 2: a) An attention map. Blue and green denotes the prompt and response tokens, respectively. b) The corresponding attention graph $G$. Prompt tokens $P$ are located on the left, response tokens $R$ — on the right. To keep figure neat, we only plot the edges with an attention score of no less than $0.15$. c) The minimum spanning forest attaching $R$ to $P$.
Through analysis of divergence patterns across different heads, we identified a subset of attention heads that consistently assign higher divergence scores to hallucinated samples (see Figure 1), revealing their implicit “awareness” of hallucinations. TOHA utilizes the average divergence values from these specific heads as hallucination scores. Crucially, these heads exhibit consistent behavior across different datasets, enabling strong transferability of our method.
Our main contributions can be summarized the following:
  • We propose TOHA, a training-free method based on the topological divergences of attention graphs. While efficient — TOHA operates up to an order of magnitude faster than methods of comparable quality and requires minimal annotated data — our method demonstrates strong in-domain performance and maintains domain transferability across different tasks.
  • The existence of hallucination-aware attention heads is discovered: calculating topological divergences from just six specific heads is enough for reliable hallucination detection, irrespective of the dataset.
  • Our experiments show TOHA consistently matches or exceeds state-of-the-art performance on all benchmarks when applied to modern open-source LLMs of varying scales (7B to 13B parameters).
2 Background
2.1 Attention matrix as a weighted graph
Modern LLMs are mainly based on the self-attention mechanism, introduced in Vaswani et al. (2017). Let $X\in\mathbb{R}^{n\times d}$ be a matrix consisting of $d$-dimensional representations of $n$ tokens, $W_{Q},\,W_{K},\,W_{V}\in\mathbb{R}^{d\times d}$ be trainable projection matrices. Given a set of queries $Q=XW_{Q}\in\mathbb{R}^{n\times d}$, a set of keys $K=XW_{K}\in\mathbb{R}^{n\times d}$, and corresponding values $V=XW_{V}\in\mathbb{R}^{n\times d}$, the attention mechanism calculates a weighted sum of the values as follows:
$\operatorname{Attention}(Q,K,V)=W(Q,K)V,$ (1)
where $W(Q,K)$ is an attention matrix
$\operatorname{W}=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right),$ (2)
and each entry $w_{ij}$ in it captures how strongly token $i$ attends to token $j$, $i\geq j$ for a decoder, with larger $w_{ij}$ indicating closer relationship.
An attention matrix $W$ can be reframed as a complete weighted graph $G$ where tokens are vertices and weights $w_{ij}$ represent the strength of connections between them. From the perspective of topological data analysis, however, it is more convenient to consider these weights as pseudo-distances rather than correlation measures. Hence, we reassign the edge weights of such a graph to equal $1-w_{ij}$, creating what we call attention graphs.
In the generation process, the vertices of such graphs naturally partition into two distinct subsets: prompt tokens $P$ and response tokens $R$ generated by the model (see Figure 2b). This split allows us to analyze the topological relationships between input and output content.
2.2 Manifold Topology Divergence
One way to compare two data manifolds, $\mathcal{M}$ and $\mathcal{N}$, approximated by point clouds $M$ and $N$ is the $\operatorname{MTop-Div}(M,N)$ topological measure Barannikov et al. (2021). This divergence is based on the $\operatorname{Cross-Barcode}(M,N)$, which is a set of intervals $\{(b_{i},d_{i})\}_{i=1}^{n}$ corresponding to “births” and “deaths” of independent topological features that distinguish the point cloud $N$ from the union $M\cup N$. The further $\operatorname{Cross-Barcode}(M,N)$ is from an empty set, the more the data manifold $\mathcal{N}$ differs from $\mathcal{M}$ in its topological structure. To measure the distance from an empty set, the sum of lengths of intervals in $\operatorname{Cross-Barcode}(M,N)$ was taken in loc. cit.:
$\operatorname{MTop-Div}(M,N)=\sum\limits_{i=1}^{n}|d_{i}-b_{i}|.$
3 Method
Given an attention matrix for the (prompt + response) text, we construct the attention graph, imitating a data manifold of the text, and study its relation with the weighted subgraph, imitating the data submanifold of the prompt. We measure the topological divergence between these graphs, assuming responses that are consistent with the context would result in lower divergence values.
3.1 $\operatorname{MTop-Div}$ for attention graphs
The $\operatorname{MTop-Divergence}(M,N)$ measure was originally developed for data manifolds, relying on metric space properties to some extent. We adapt this concept to quantify divergence between attention graphs and their subgraphs, where traditional metric axioms do not apply, while preserving the measure’s comparative utility.
Let $R$ and $P$ be the sets of response and the prompt vertices in the attention graph $G$. We set to zero the edge weights between the $P$ vertices, denote $w_{(R\cup P)/P}$ the resulting matrix of edge weights, and define $\operatorname{Cross-Barcode}_{i}(R,P)$ as the $i$-th homology barcode of the Vietoris-Rips simplicial complex $VR_{\alpha}(G,w_{(R\cup P)/P})$. We define $\operatorname{MTop-Div}(R,P)$ as the total sum of interval lengths in $\operatorname{Cross-Barcode}_{0}(R,P)$, where we consider the 00-dimensional homology group $H_{0}$. In this context, we can prove that this score is equivalent to the total edge length of the minimum spanning forest (MSF) connecting $R$ to $P$.
Properties of MTop-Div for attention graphs.
Here, we only list the properties relevant to hallucination detection; for more properties and proofs, see Appendix A.
Proposition 3.1.
The following holds for any attention graph $G$ with vertex set $V_{G}$ and its complementary vertex subsets $P,R$, where $P\cup R=V_{G}$ and $P\cap R=\varnothing$.
1. (Formula.) $\operatorname{MTop-Div}(R,P)$ value equals the length of the MSF attaching $R$ to $P$.
2. (Stability.) If the weights of $G$ change by no more than $\varepsilon$, then the corresponding $\operatorname{MTop-Div}(R,P)$ changes by no more than $\delta=\varepsilon|R|$.
3. (Connection with hallucinations.) The normalized divergence value $\frac{1}{|R|}\operatorname{MTop-Div}(R,P)=0$ iff the MSF attaches every response token to a prompt token by a subtree with attention weights $=1$.
The stability property guarantees that similar attention patterns yield similar hallucination scores, making the metric’s behavior consistent and predictable. The latter captures the intuitive relationship between divergence and response quality: well-grounded responses (closely tied to the context) produce small divergence values, while hallucinations (occurring when evidence is missing) lead to smaller attention weights and consequently larger divergence values. Together, these properties enable the metric to reliably measure how strongly a response connects to its context while remaining robust to minor attention variations.
3.2 Hallucination-aware heads
We hypothesize, inspired by prior investigations in LLM interpretability Voita et al. (2019); Gould et al. (2024), that particular attention heads exhibit distinct patterns related to hallucinations. To identify such heads, we analyzed head-specific topological divergences as follows.
Denote by $h_{ij}$ the $j$-th attention head from the layer $i$. For the specific data sample $s$ and head $h_{ij}$, let $G^{s}_{ij}$ be the corresponding attention graph, $P^{s}_{ij},R^{s}_{ij}$ — its prompt and response vertex subsets.
We examined typical values of the average distance between hallucinated and grounded training examples for different heads and layers:
$\Delta_{ij}=\frac{1}{|S_{\mathrm{hallu}}|}\sum\limits_{s\in S_{\mathrm{hallu}}}d_{ij}(s)-\frac{1}{|S_{\mathrm{gr}}|}\sum\limits_{s\in S_{\mathrm{gr}}}d_{ij}(s),$
where $S_{\mathrm{hallu}}$ stands for all hallucinated samples from the training set, $S_{\mathrm{gr}}$ stands for all grounded training samples, and
$d_{ij}(s)=\frac{1}{|R_{ij}^{s}|}\operatorname{MTop-Div}(R^{s}_{ij},P^{s}_{ij}).$
The sample differences obtained for three datasets are displayed in Figure 3. Each dot represents an individual attention head, with its $x$-coordinate indicating its $\Delta_{ij}$ value on dataset (A) and its $y$-coordinate the corresponding value on dataset (B). For each dataset pair, we highlight the top three most separating heads for dataset (A) in pink. Notably, these heads consistently appear in the upper-right corner of the plot, indicating that they also exhibit strong separation on dataset (B). This observation suggests that these heads are inherently attuned to hallucination patterns, regardless of the dataset.
Figure 3: $\Delta_{ij}$ values for $ij$-th heads. Vertical axis corresponds to the difference on dataset (B), horizontal — to the one on dataset (A). The heads that separate samples best are highlighted in pink. Model names for a row are on the left side, datasets: MS MARCO, CNN/DM + Recent News, CoQA.
3.3 TOHA
The existence of universal hallucination patterns in the attention heads underlies our efficient method TOHA, detailed in Algorithm 1. It uses two small, annotated probe sets, $S_{h}$ (containing hallucinated samples) and $S_{g}$ (containing grounded samples), to rank model heads by their separation capability based on their $\Delta_{ij}$ values, where $ij$ denotes a head index, and select the most relevant ones. In our experiments, the combined size of the probe sets is kept small (see Figure 4 for the number of the required samples analysis). During testing, hallucination scores are computed as the average topological divergence from the top $N_{\mathrm{opt}}$ heads, where $N_{\mathrm{opt}}$ is a hyperparameter tuned on the set $V=S_{h}\cup S_{g}$. For computational efficiency, we limit $N_{\mathrm{opt}}$ to a maximum of 6 in all experiments.
Algorithm 1 TOHA algorithm
$d_{ij}(s)$ — topological divergences for samples;
$S_{h},\,S_{g}$ — probe sets;
$V=S_{h}\cup S_{g}$ — validation set;
$T$ — test set;
$N_{\mathrm{max}}$ — max number of selected heads.
procedure TOHA heads selection
     $\Delta_{ij}\leftarrow\frac{1}{|S_{h}|}\sum\limits_{s\in S_{h}}d_{ij}(s)-\frac{1}{|S_{g}|}\sum\limits_{s\in S_{g}}d_{ij}(s)$
     $H\leftarrow\mathrm{sort\_descending}(h_{ij},\text{key}=\Delta_{ij})$
     $N,N_{\mathrm{opt}}\leftarrow 1,1$
     $H_{\mathrm{subset}}\leftarrow\varnothing$triangleright Optimal heads set.
     $\text{AUROC}_{\max}\leftarrow 0$
     $p_{s}=0,s\in V$triangleright Hallucination scores.
     while $N\leq N_{\max}$ do
         $H_{\mathrm{subset}}\leftarrow H_{\mathrm{subset}}\cup\{h_{N}\}$
         for $s\in V$ do
              $p_{s}\leftarrow\frac{N-1}{N}p_{s}+\frac{1}{N}d_{h_{N}}(s)$
         end for
         $\text{auroc}\leftarrow\text{AUROC}(\{y_{s}\}_{s\in V},\{p_{s}\}_{s\in V})$
         if $\text{auroc}>\text{AUROC}_{\max}$ then
              $\text{AUROC}_{\max}\leftarrow\text{auroc}$
              $N_{\mathrm{opt}}\leftarrow N$
         end if
         $N\leftarrow N+1$
     end while
end procedure
procedure TOHA prediction
     for $s\in T$ do $\triangleright$ Prediction on the test set.
         $p_{s}\leftarrow\frac{1}{N_{\mathrm{opt}}}\sum\limits_{i=1}^{N_{\mathrm{opt}}}d_{h_{i}}(s)$
     end for
end procedure
4 Experiments
Datasets.
We evaluated our approach on four datasets: RAGTruth Niu et al. (2023) (we considered its’ two separate benchmarks: QA based on MS MARCO Nguyen et al. (2016) and summarization based on CNN/DM Nallapati et al. (2016)) combined with news articles from an unnamed news platform), CoQA Reddy et al. (2019), SQuAD Rajpurkar et al. (2016), and XSum Narayan et al. (2018). The RAGTruth dataset consists of manually annotated responses of several LLMs in the RAG setting. The annotations are word-level; we, in turn, predict response-level labels, considering a response hallucinated if it contains at least one hallucination span. For the three latter datasets, we sampled LLM responses and annotated them automatically with GPT-4o Hurst et al. (2024). Consistency with Human-GPT-4o label (Table 7) validated this approach, matching prior work Bavaresco et al. (2024); see Appendix C for more details.
Models.
We used five popular open-source LLMs: LLaMA-2-7B-chat, LLaMA-2-13B-chat, LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct-v0.1, and Qwen2.5-7B-Instruct. Note that the RAGTruth dataset does not contain responses for LLaMA-3.1-8B and Qwen-2.5-7B; therefore, we only conducted experiments on SQuAD, CoQA, and XSum for these models.
Baselines.
We compare TOHA with seven baselines: perplexity Ren et al. (2023), max entropy Fadeeva et al. (2024), Haloscope Du et al. (2024), LLM-Check Sriramanan et al. (2024), semantic entropy Farquhar et al. (2024), EigenScore Chen et al. (2024), and SelfCheckGPT Manakul et al. (2024). Appendix E provides information on implementation details.
Table 1: ROC AUC ()($\uparrow$)( ↑ ) of hallucination detection techniques for three LLMs. The best results for each model are highlighted in bold, and the second best are underlined.
Method Single MS MARCO CNN/DM + CoQA SQuAD XSum
generation Recent News
Mistral-7B
SelfCheckGPT [1] 0.62 $\pm$ 0.03 0.63 $\pm$ 0.04 0.93 $\pm$ 0.02 0.82 $\pm$ 0.03 0.71 $\pm$ 0.03
Semantic entropy [2] 0.54 $\pm$ 0.06 0.57 $\pm$ 0.04 0.84 $\pm$ 0.02 0.71 $\pm$ 0.05 0.65 $\pm$ 0.05
EigenScore [3] 0.58 $\pm$ 0.02 0.55 $\pm$ 0.04 0.73 $\pm$ 0.02 0.54 $\pm$ 0.04 0.57 $\pm$ 0.03
Haloscope [4] 0.56 $\pm$ 0.05 0.54 $\pm$ 0.05 0.84 $\pm$ 0.02 0.97 $\pm$ 0.03 0.59 $\pm$ 0.05
LLM-Check [5] 0.5 $\pm$ 0.06 0.57 $\pm$ 0.03 0.62 $\pm$ 0.04 0.56 $\pm$ 0.06 0.57 $\pm$ 0.03
Perplexity [6] 0.67 $\pm$ 0.03 0.62 $\pm$ 0.04 0.77 $\pm$ 0.01 0.45 $\pm$ 0.06 0.66 $\pm$ 0.03
Max entropy [7] 0.66 $\pm$ 0.05 0.6 $\pm$ 0.05 0.73 $\pm$ 0.02 0.73 $\pm$ 0.04 0.71 $\pm$ 0.03
TOHA (ours) 0.76 $\pm$ 0.04 0.63 $\pm$ 0.04 0.89 $\pm$ 0.02 0.84 $\pm$ 0.01 0.61 $\pm$ 0.03
LLama-2-7B
SelfCheckGPT [1] 0.6 $\pm$ 0.02 0.62 $\pm$ 0.05 0.78 $\pm$ 0.04 0.59 $\pm$ 0.05 0.70 $\pm$ 0.06
Semantic entropy [2] 0.54 $\pm$ 0.05 0.54 $\pm$ 0.04 0.76 $\pm$ 0.03 0.63 $\pm$ 0.06 0.63 $\pm$ 0.05
EigenScore [3] 0.54 $\pm$ 0.02 0.52 $\pm$ 0.04 0.68 $\pm$ 0.03 0.53 $\pm$ 0.05 0.63 $\pm$ 0.06
Haloscope [4] 0.53 $\pm$ 0.06 0.48 $\pm$ 0.02 0.74 $\pm$ 0.04 0.58 $\pm$ 0.04 0.58 $\pm$ 0.08
LLM-Check [5] 0.46 $\pm$ 0.02 0.49 $\pm$ 0.03 0.6 $\pm$ 0.03 0.58 $\pm$ 0.05 0.58 $\pm$ 0.08
Perplexity [6] 0.66 $\pm$ 0.01 0.57 $\pm$ 0.04 0.73 $\pm$ 0.03 0.57 $\pm$ 0.09 0.58 $\pm$ 0.05
Max entropy [7] 0.66 $\pm$ 0.03 0.57 $\pm$ 0.05 0.71 $\pm$ 0.04 0.64 $\pm$ 0.02 0.56 $\pm$ 0.05
TOHA (ours) 0.67 $\pm$ 0.02 0.57 $\pm$ 0.05 0.88 $\pm$ 0.04 0.9 $\pm$ 0.04 0.63 $\pm$ 0.02
LLaMA-2-13B
SelfCheckGPT [1] 0.57 $\pm$ 0.04 0.60 $\pm$ 0.04 0.86 $\pm$ 0.04 0.78 $\pm$ 0.03 0.61 $\pm$ 0.05
Semantic entropy [2] 0.61 $\pm$ 0.03 0.51 $\pm$ 0.05 0.75 $\pm$ 0.05 0.72 $\pm$ 0.02 0.64 $\pm$ 0.04
EigenScore [3] 0.56 $\pm$ 0.04 0.49 $\pm$ 0.04 0.45 $\pm$ 0.04 0.51 $\pm$ 0.04 0.55 $\pm$ 0.02
Haloscope [4] 0.55 $\pm$ 0.05 0.50 $\pm$ 0.01 0.65 $\pm$ 0.04 0.54 $\pm$ 0.03 0.59 $\pm$ 0.01
LLM-Check [5] 0.44 $\pm$ 0.04 0.57 $\pm$ 0.04 0.54 $\pm$ 0.03 0.53 $\pm$ 0.05 0.63 $\pm$ 0.04
Perplexity [6] 0.65 $\pm$ 0.02 0.59 $\pm$ 0.02 0.58 $\pm$ 0.04 0.5 $\pm$ 0.05 0.58 $\pm$ 0.05
Max entropy [7] 0.58 $\pm$ 0.03 0.55 $\pm$ 0.06 0.70 $\pm$ 0.06 0.78 $\pm$ 0.04 0.56 $\pm$ 0.04
TOHA (ours) 0.69 $\pm$ 0.01 0.54 $\pm$ 0.04 0.93 $\pm$ 0.03 0.92 $\pm$ 0.02 0.66 $\pm$ 0.05
Table 2: ROC AUC ()($\uparrow$)( ↑ ) of hallucination detection techniques. The best results for each model are highlighted in bold, and the second best are underlined.
Method Single gen. SQuAD CoQA XSum
LLaMA-3.1-8B
SelfCheckGPT [1] 0.79 $\pm$ 0.05 0.76 $\pm$ 0.07 0.81 $\pm$ 0.02
Semantic entropy [2] 0.58 $\pm$ 0.04 0.83 $\pm$ 0.05 0.49 $\pm$ 0.05
EigenScore [3] 0.52 $\pm$ 0.07 0.82 $\pm$ 0.06 0.49 $\pm$ 0.05
Haloscope [4] 0.85 $\pm$ 0.02 0.56 $\pm$ 0.07 0.55 $\pm$ 0.04
LLM-Check [5] 0.48 $\pm$ 0.04 0.5 $\pm$ 0.08 0.56 $\pm$ 0.03
Perplexity [6] 0.82 $\pm$ 0.02 0.69 $\pm$ 0.04 0.61 $\pm$ 0.04
Max entropy [7] 0.5$\pm$ 0.04 0.53 $\pm$ 0.05 0.47 $\pm$ 0.01
TOHA (ours) 0.85 $\pm$ 0.02 0.73 $\pm$ 0.01 0.63 $\pm$ 0.04
Qwen2.5-7B
SelfCheckGPT [1] 0.62 $\pm$ 0.06 0.84 $\pm$ 0.04 0.74 $\pm$ 0.04
Semantic entropy [2] 0.68 $\pm$ 0.03 0.77 $\pm$ 0.06 0.65 $\pm$ 0.04
EigenScore [3] 0.67 $\pm$ 0.04 0.66 $\pm$ 0.09 0.51 $\pm$ 0.04
Haloscope [4] 0.59 $\pm$ 0.05 0.71 $\pm$ 0.05 0.58 $\pm$ 0.05
LLM-Check [5] 0.5 $\pm$ 0.05 0.54 $\pm$ 0.1 0.55 $\pm$ 0.05
Perplexity [6] 0.63 $\pm$ 0.06 0.67 $\pm$ 0.08 0.65 $\pm$ 0.05
Max entropy [7] 0.73 $\pm$ 0.05 0.74 $\pm$ 0.05 0.52 $\pm$ 0.08
TOHA (ours) 0.8 $\pm$ 0.02 0.69 $\pm$ 0.05 0.69 $\pm$ 0.02
Main results.
The results of our experiments are provided in Tables 12. We evaluate TOHA against state-of-the-art hallucination detection methods and demonstrate its competitive performance, consistently securing first or second place across most benchmark datasets. TOHA significantly outperforms uncertainty-based baselines and matches the accuracy of consistency-based approaches, achieving a notable $13.4\%$ improvement on the challenging MS MARCO dataset that includes long and detailed model responses. While SelfCheckGPT emerges as TOHA’s closest competitor, it relies on additional generations, incurring substantial computational overhead. Comparison of the best-performing baselines (Figure 6) confirms that TOHA reduces inference time by an order of magnitude compared to methods of similar quality.
Compared to Haloscope Du et al. (2024), which operates with limited annotated data and large-scale unlabeled data, TOHA not only delivers superior performance but also eliminates the need for unannotated generations for hyperparameter tuning and a separate classifier training. This makes TOHA less data-dependent and more practical for real-world applications.
Another interesting comparison is with LLM-Check Sriramanan et al. (2024), which also uses attention maps to compute hallucination scores—specifically, by averaging the log determinant of attention maps from a single pre-selected layer. However, our TOHA achieves superior performance, demonstrating that not all attention heads contribute equally to hallucination detection. By employing a topology-based head selection strategy, we significantly enhance detection quality.
Generalizability to different data distributions.
From a deployment perspective, hallucination detection methods must remain robust to shifts in input data distribution, given the inherent diversity of real-world user queries. To evaluate TOHA’s robustness in this regard, we conducted transfer experiments on Mistral-7B (see Figure 4a). The results highlight TOHA’s strong transferability: for the XSum and CNN/DM datasets, performance changes in transfer settings fall within the standard deviation. For the remaining datasets, TOHA maintains competitive performance compared to baseline methods (Table 1), demonstrating its adaptability to diverse data distributions.
How large should the probe sets be?
As previously mentioned, TOHA requires only a small set of samples to identify ”hallucination-aware" attention heads. To assess its sensitivity to probe set size, we conducted an ablation study (Figure 4). The results demonstrate TOHA’s robustness to limited annotated data: even with just $50$ samples, performance does not drop significantly and mostly remains stable as the probe set size increases.
Figure 4: (a) Generalizability between the datasets, model: Mistral-7B-Instruct. The vertical axis corresponds to the origin of the probe set, the horizontal axis  to the test dataset. (b)-(c): Detection quality dependence on the size of a probe set, models: Mistral-7B (left), LLama-2-7B (right).
What do hallucination patterns look like?
As detailed in Section 3, the topological divergences we employ characterize the MSF connecting the vertices $R$ of the response to the vertices $P$ of the prompt. For hallucination-aware heads, we analyzed MSF patterns distinguishing hallucinated and grounded samples. A key finding is that hallucinated samples frequently exhibit strong attention to the <s> token, whereas grounded samples tend to attend to <s> less (Figure 5).
To verify the significance of <s> in hallucination detection, we conducted an ablation experiment: after removing <s> from the texts, we recomputed the TOHA hallucination scores for the selected hallucination-aware heads. The results (Table 3) show a significant performance drop, confirming that attention to <s> is a critical indicator of hallucination. This finding aligns with prior work demonstrating the influential role of <s> in LLM mechanisms Barbero et al. (2025). However, using the average attention to <s> alone as a hallucination score proves insufficient (see Table 9 in Appendix D). In contrast, our proposed score, which incorporates the intricate structure of attention maps, demonstrates far greater discriminative power for this task.
Table 3: TOHA performance with and without <s> token. TOP-1 results are highlighted with bold font.
Dataset with <s> w/o <s>
Mistral-7B
MS MARCO 0.67 $\pm$ 0.02 0.56 $\pm$ 0.02
CoQA 0.88 $\pm$ 0.02 0.32 $\pm$ 0.03
LLaMA-2-7B
MS MARCO 0.76 $\pm$ 0.04 0.66 $\pm$ 0.03
CoQA 0.89 $\pm$ 0.02 0.56 $\pm$ 0.04
Figure 5: Attention to <s>: a) hallucinated sample and b) grounded one. Green color denotes edges and nodes that correspond to grounded tokens of a generation, yellow color — hallucinated ones. Model: Mistral-7B.
Figure 6: Comparison of methods’ inference time in seconds. The measurements were obtained for $16$ random samples from MS MARCO, model: Mistral-7B. For the SelfCheckGPT, $20$ additional answers were generated for each sample.
5 Related works
Hallucination detection methods.
The problem of hallucinations in LLMs has attracted significant attention recently Zhang et al. (2023); Huang et al. (2023); Wang et al. (2024). Consistency-based methods  Manakul et al. (2024); Chen et al. (2024); Kuhn et al. (2023); Qiu and Miikkulainen (2024); Nikitin et al. (2024) that use the diversity of multiple LLM responses as a hallucination score offer robust detection but impose significant computational overhead. Surface-level techniques like perplexity and logit entropy Fadeeva et al. (2024); Malinin and Gales (2021) analyze model confidence directly from output distributions — efficient but limited in detection capability as they neglect the model’s rich internal representations. Hidden states-based classifiers Azaria and Mitchell (2023); Sky et al. (2024); Zhou et al. (2025) require extensive annotated datasets, which are scarce in the public domain Zhang et al. (2023). This issue was partially addressed by the Haloscope Du et al. (2024), which leverages unlabeled data “in the wild” with minimum annotated data needed for hyperparameter selection. Attention map-based methods represent a promising yet underdeveloped direction. Current techniques either rely on large labeled data, e.g., Lookback Lens Chuang et al. (2024), or exploit only simple attention graph properties, such as self-loop weights in LLM-Check Sriramanan et al. (2024). This leaves a critical research gap: training-free methods that fully leverage the rich structural information encoded in attention relationships remain underexplored.
Topological Data Analysis (TDA) in NLP.
Topological Data Analysis (TDA) is a mathematical framework for extracting multi-scale structural patterns from data using principles from topology and computational geometry Chazal and Michel (2017); Hensel et al. (2021). Recent years have seen growing interest in applying TDA to natural language processing (NLP) tasks to study textual structural properties Uchendu and Le (2024). For example,  Tulchinskii et al. (2024) leveraged persistent homology to estimate the intrinsic dimensionality of CLS embeddings for detecting machine-generated text. Other work has demonstrated the utility of topological features derived from transformer attention matrices — treated as weighted graphs — for diverse NLP applications. These include uncertainty quantification Kostenok et al. (2023) and grammatical acceptability classification Proskurina et al. (2023b), where topological features extracted from the attention graphs were used as input to train auxiliary classifiers.
6 Conclusion
This paper introduces TOHA, a novel hallucination detection method based on the topological divergence of attention maps. At its core, TOHA leverages our key observation that specific attention heads exhibit consistent patterns during hallucinations — regardless of the dataset. TOHA computes hallucination scores by averaging the topological divergences from these heads, and we formally prove several stability properties to ensure these scores are reliable. Additionally, we explored the behaviour of “hallucination-aware” heads, discovering that the attention to <s> token plays an important role in their discriminative ability. This importance of <s> aligns well with prior work Barbero et al. (2025).
Extensive experiments show that TOHA is a robust alternative to existing approaches, matching or surpassing state-of-the-art baselines like SelfCheckGPT Manakul et al. (2024). Notably, TOHA is significantly more efficient, operating up to an order of magnitude faster than methods of comparable quality. We further validate TOHA’s transferability, demonstrating its robustness to shifts in data distribution — a critical advantage for real-world deployment, where LLM inputs are far more diverse and complex than benchmark examples.
In summary, TOHA delivers state-of-the-art detection performance while combining efficiency and solid generalizability, making it particularly suited for practical applications.
Limitations
While TOHA demonstrates strong performance and efficiency, several limitations warrant discussion.
RAG scenario. While TOHA operates effectively in the RAG scenario under the assumption that the provided context contains the correct answer, we recognize that this condition may not always hold in real-world applications. This limitation points to an important direction for future research, where the method could be extended to handle cases of incomplete or unreliable context knowledge.
Model-specific dependencies. TOHA’s effectiveness relies on identifying “hallucination-aware” attention heads, which may vary across LLM architectures. While our experiments cover popular open-source models (e.g., LLaMA, Mistral), further validation is needed for proprietary or larger models (e.g., GPT-4, Claude).
Multimodal extensions. The current framework operates solely on text. Adapting TOHA to multimodal settings (e.g., vision-language models) would require redefining attention graphs across heterogeneous data modalities.
Appendix A Topological data analysis: background
A simplicial complex $S$ is a collection of simplices such that every face of a simplex $\sigma\in S$ is also in $S$. Simplices are the higher-dimensional generalizations of triangles; a 0-simplex is a vertex, a 1-simplex is an edge, a 2-simplex is a triangle, and so forth. Formally, given a finite set $X$, an $n$-simplex $\sigma$ is an $(n+1)$ subset of $X$. Simplicial complexes are fundamental objects in algebraic and combinatorial topology, serving as a discrete analog to topological spaces.
The Vietoris-Rips complex $VR_{\varepsilon}(X)$ of a weighted graph $G=(V_{G},E_{G})$ with distance threshold $\varepsilon>0$ is defined as follows:
$\mathrm{VR}_{\varepsilon}(G)=\left\{\sigma\subseteq V_{G}\ \bigg{|}\ \forall\,v_{i},v_{j}\in\sigma,\ w(e_{ij})\leq\varepsilon\right\},$
where $w$ is the edge weight function associated with $G$.
Homology groups $H_{k}$ are invariants used in algebraic topology to study the topological properties of a space. Let $C_{k}(S)$ denote vector space over $\mathbb{Z}/2\mathbb{Z}$, with the basis consisting of $k$-dimensional simplices of $S$. Elements of $C_{k}$ are called chains. Formally, homology groups are derived from a chain complex $(C_{\bullet},\partial_{\bullet})$, which is a sequence of $C_{k}$ connected by boundary maps $\partial_{k}$:
$C_{\bullet}:\cdots\rightarrow C_{k+1}\xrightarrow{\partial_{k+1}}C_{k}\xrightarrow{\partial_{k}}\cdots,\,$
$\partial_{k}\circ\partial_{k+1}=0.$
The $k$-th homology group $H_{k}$ is defined as the quotient of the group of $k$-cycles (chains whose boundary is zero) by the group of $k$-boundaries (chains that are the boundary of a $(k+1)$-chain). Mathematically, this is expressed as:
$H_{k}(S)=Z_{k}(S)/B_{k}(S),$
where $Z_{k}=\mathrm{ker}\,\partial_{k}=\{c\in C_{k}\,|\,\partial_{k}(c)=0\}$ and $B_{k}=\mathrm{im}\,\partial_{k+1}=\{\partial_{k+1}(c)\,|\,c\in C_{k+1}\}$ is the group of $k$-boundaries. The elements of $H_{k}(S)$ represent various $k$-dimensional topological features in $S$. Elements of a basis in $H_{k}(S)$ correspond to a set of basic topological features.
A filtration of simplicial complexes $\mathcal{F}$ is a family of nested simplicial complexes:
$\mathcal{F}:\varnothing\subseteq S_{1}\subseteq S_{2}\subseteq\dots\subseteq S_{n}=S,$
where each $S_{k}$ is a simplicial complex itself. In practice, the filtrations of simplicial complexes are usually obtained for sequences of increasing thresholds $0<\varepsilon_{1}<\dots<\varepsilon_{n}$. For example, simplicial complexes $VR_{\varepsilon_{i}}(X)$ form a filtration
$\mathcal{F}_{VR}(X):$ $\varnothing\subseteq VR_{\varepsilon_{1}}(X)\subseteq VR_{\varepsilon_{2}}(X)\subseteq\dots$
$\subseteq VR_{\varepsilon_{n}}(X)=VR(X).$
[…]
Figure 7: $H_{0}$ barcode construction. As the threshold increases, the separate connected components merge, resulting in the death of topological features. The horizontal axis is a sequence of thresholds $\varepsilon$, and each horizontal bar corresponds to a single feature.
Appendix B MTop-Div on graphs properties
Basic properties of MTop-Div for attention graphs.
Now we consider specific properties for our adaptation of $\operatorname{MTop-Div}(R,P)$.
Proposition B.1.
The following holds for any attention graph $G$ and its complementary vertex subsets $P,R\subset~{}V_{G}$.
  • $\operatorname{MTop-Div}(R,P)$ value equals the length of the MSF attaching $R$ to $P$.
  • Let the natural norm on the cross-barcodes be defined as follows:
    $\lVert\operatorname{Cross-Barcode}_{0}\rVert_{B}=\\ \max_{[b_{j},d_{j}]\in\operatorname{Cross-Barcode}_{0}}(d_{j}-b_{j}).$ (3)
    The norm of $\operatorname{Cross-Barcode}_{0}(R,P)$ lays in the interval [0, 1]01[0,\,1][ 0 , 1 ]:
    $0\leq\lVert\operatorname{Cross-Barcode}_{0}(R,P)\rVert_{B}\leq 1$ (4)
  • The divergence itself is bounded by
    $0\leq\operatorname{MTop-Div}(R,P)\leq|R|.$ (5)
The second and third statements are immediately obtained from the properties of an attention matrix: all its weights lie between 00 and $1$.
The following property formalizes the intuition behind our metric — it measures the strength of the response’s connection to the prompt through multiscale topological features of the attention graph.
Proposition B.2.
(Exact sequence.) For any $\alpha$, the following sequence of natural maps of homology groups is exact
$(\mathbb{Z}/2\mathbb{Z})^{\lvert P\rvert}$ $\xrightarrow{r_{2}}H_{0}(VR_{\alpha}(G))\xrightarrow{r_{1}}$
$\xrightarrow{r_{1}}H_{0}(VR_{\alpha}(G,w_{(R\cup P)/P}))\xrightarrow{r_{0}}0.$
Proof of Proposition B.2.
We have to check the definition of the exact sequence: $\text{Ker}(r_{i})=\text{Im}(r_{i+1})$. For a pair $r_{0},\,r_{1}$, it is equivalent to the surjectvity of $r_{1}$. The $H_{0}$ homology group of a graph corresponds to the connected components of the graph. The set of edges $E_{(G,w)}^{\leq\alpha}=\{e\in E_{G}|w_{e}\leq\alpha\}$ is always a subset in the analogous set of the weighted graph $(G,w_{(R\cup P)/P})$ with all weight edges between $P$ vertices set to zero. Therefore, the map $r_{1}$ between their connected components is surjective. Similarly, the kernel of the map $r_{1}$ is spanned by the differences of two connected components, which are merged after adding some of the edges between $P$ vertices, and any such difference lies in the image of the map $r_{2}$. Also, any two vertices from $P$ belong to the same connected component in the graph $(G,w_{(R\cup P)/P}\leq\alpha)$, hence the image of $r_{2}$ is in the kernel of $r_{1}$. Therefore, the considered sequence is exact indeed. ∎
Proof of Proposition 3.1.
1. The $0-$th Cross-Barcode coincides with the set of edges in the minimal spanning tree of the weighted graph $G$ with all the weights within $P$ vertex subset equal zero. Excluding the zero weight edges, this edge set coincides with the minimal spanning forest attaching the vertex set $R$ to $P$ vertices. ∎
2. Denote by $\operatorname{MSF}(R,P)$ the minimum spanning forest attaching $R$ to $P$. Note that we have properties B.1, so
$\operatorname{MTop-Div}(R,P)=\sum_{e\in\operatorname{MSF}(R,P)}w(e).$ (6)
Therefore, we have to show that the weight of $\operatorname{MSF}(R,P)$ does not change significantly when all weights are changed by no more than $\varepsilon$.
There are two possibilities: 1) after a change, all MSF edges remain the same, or 2) some edges are replaced with other edges. In the first case, it is obvious that the total sum of edge weights changes by no more than $\delta=\varepsilon\cdot\#\text{edges}(\operatorname{MSF}(R,P))=\varepsilon\cdot|R|$. Consider the second case. Denote by $\operatorname{MSF}_{\mathrm{prev}}$ the original MSF, by $\operatorname{MSF}_{\mathrm{new}}$ — the MSF after the change; let $w$ be the edge weight function before the change, $\hat{w}$ — after the change. The following inequalities hold:
$\hat{w}(\operatorname{MSF}_{\mathrm{new}})$ $<\hat{w}(\operatorname{MSF}_{\mathrm{prev}});$ (7)
$w(\operatorname{MSF}_{\mathrm{prev}})-\delta\leq\hat{w}$ $(\operatorname{MSF}_{\mathrm{prev}})\leq$
$\leq w(\operatorname{MSF}_{\mathrm{prev}})+\delta;$ (8)
$w(\operatorname{MSF}_{\mathrm{new}})-\delta\leq\hat{w}$ $(\operatorname{MSF}_{\mathrm{new}})\leq$
$\leq w(\operatorname{MSF}_{\mathrm{new}})+\delta;$ (9)
$w(\operatorname{MSF}_{\mathrm{new}})$ $\geq w(\operatorname{MSF}_{\mathrm{prev}}).$ (10)
From (7)-(8) follows that $\hat{w}(\operatorname{MSF}_{\mathrm{new}})<w(\operatorname{MSF}_{\mathrm{prev}})+\delta$; from (9)-(10) follows that $\hat{w}(\operatorname{MSF}_{\mathrm{new}})\geq w(\operatorname{MSF}_{\mathrm{prev}})-\delta$. ∎
3. Follows obviously from the MSF formula for $\operatorname{MTop-Div}(R,P)$ and attention map properties.
Appendix C Datasets
SQuAD Rajpurkar et al. (2016) and CoQA Reddy et al. (2019) are widely used English question-answering benchmarks that have facilitated the development of hallucination detection datasets Kuhn et al. (2023); Manakul et al. (2024). Similarly, XSum Narayan et al. (2018), a dataset of news articles with one-sentence summaries, is commonly employed in hallucination detection research for abstractive summarization Shi et al. (2024); Cao et al. (2022). To assess LLM performance, we used GPT-4o to annotate responses to questions sourced from SQuAD, CoQA, and summarization tasks from XSum.
SQuAD CoQA
Given the context, answer the question in a brief but complete sentence. Once upon a time, in a quiet village, there lived a kind old baker named Henry.
Note that your answer should be strictly based on the given context. He was known for his delicious bread and warm smile. One day, a traveler arrived,
In case the context does not contain the necessary information to answer the question, tired and hungry, Henry welcomed him with a fresh loaf.
please reply with “Unable to answer based on given context”. Q: What was Henry known for?
Context: A: Baking delicious bread.
Once upon a time, in a quiet village, there lived a kind old baker named Henry. Q: What else?
He was known for his delicious bread and warm smile. One day, a traveler arrived, A: Warm smile.
tired and hungry, and Henry welcomed him with a fresh loaf. Q: How did the traveler feel when he arrived?
Question: Who was known for baking delicious bread? A: Tired and hungry.
Answer: Q: What did Henry give the traveler?
Table 4: Examples of prompts used during generation for CoQA and SQuAD (we add additional delimiter spaces and formatting not present in actual prompts for better readability). SQuAD contains instructions followed by context and questions. In CoQA, the prompt has only a contextual passage followed by a question-and-answer series, with the last question being the actual one.
XSum
Please annotate potentially hallucinated model-generated summaries in the following settings.
I will provide a reference text and a model-generated summary of this text. You will judge whether the given model-generated
summary contains hallucinations. Answer "Yes" if the summary contains hallucinations, "No" if it does not, and "N/A" if you cannot decide.
Do NOT give any extra explanations.
Table 5: The prompt used during generation for the XSum dataset (we add additional delimiter spaces and formatting not present in actual prompts for better readability).
You are an AI assistant specialized in detecting hallucinations in question-answering tasks.
Your job is to analyze the given context, question, and generated answer to identify
whether the answer contains any hallucinations. Examples:
Example 1.
Context:
The city of Paris is the capital of France. It is known for its iconic landmarks
like the Eiffel Tower and Notre Dame Cathedral.
The city is situated in the northern part of the country, near the Seine River.
Question: Is Paris the capital of Germany?
Generated answer: Yes, Paris is the capital of Germany.
Hallucination: Yes.
Example 2.
Context:
The city of Paris is the capital of France.
It is known for its iconic landmarks like the Eiffel Tower and Notre Dame Cathedral.
The city is situated in the northern part of the country, near the Seine River.
Question: Is Paris the capital of Germany?
Generated answer: No, Paris is not the capital of Germany. According to the context,
Paris is the capital of France.
Hallucination: No.
You should determine if the answer contains hallucinations according to the hallucination types above.
If you cannot decide if the generated answer is a hallucination, write “N/A.” as the answer.
The answer you give MUST be ONLY “Yes.”, “No.” or “N/A.”; do NOT give ANY explanation.
Table 6: Example of annotation prompt passed to GPT-4o (we add additional delimiter spaces and formatting not present in actual prompts for better readability).
Prompt number 1 2 3 4 5
CoQA Accuracy ($\uparrow$) 0.809 ± 0.017 0.861 ± 0.015 0.742 ± 0.003 0.795 ± 0.009 0.831 ± 0.025
Precision ($\uparrow$) 0.849 ± 0.021 0.911 ± 0.007 0.771 ± 0.003 0.828 ± 0.011 0.860 ± 0.012
Recall ($\uparrow$) 0.871 ± 0.004 0.877 ± 0.019 0.877 ± 0.013 0.877 ± 0.005 0.893 ± 0.027
SQuAD Accuracy ($\uparrow$) 0.831 ± 0.003 0.857 ± 0.018 0.857 ± 0.008 0.872 ± 0.003 0.854 ± 0.007
Precision ($\uparrow$) 0.813 ± 0.002 0.831 ± 0.028 0.845 ± 0.021 0.850 ± 0.011 0.847 ± 0.007
Recall ($\uparrow$) 0.796 ± 0.008 0.839 ± 0.010 0.823 ± 0.023 0.858 ± 0.018 0.813 ± 0.017
Average Accuracy ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$)
CoQA 0.808 0.844 0.879
SQuAD 0.854 0.837 0.826
Table 7: Classification metrics of GPT-4o annotation for CoQA and SQuAD with human labels considered actual annotation. The top table shows metric scores for different variants of prompts used. The bottom table shows the metric scores averaged across all prompt variants.
C.1 Data Generation & Annotation
Generation. We generate responses from a language model (LLM) for the considered datasets, employing different prompting strategies for each dataset while keeping these strategies consistent across models (see prompt examples in Table 4). For SQuAD and XSum, responses are generated using a zero-shot approach. In contrast, for CoQA, we create queries in a few-shot manner without providing specific instructions, following Lin et al. (2024): each sample consists of a passage and a series of question-answer pairs, concluding with a final question that the model is expected to answer.
Annotation: automated vs human. We treat hallucination detection as a binary classification problem; our target indicates whether a hallucination is present anywhere in the model’s response. Two approaches to annotating model generations were considered: 1) automated annotation using an LLM (in our case, GPT-4o), and 2) manual annotation by human experts.
During the automated annotation process, we provide an LLM’s output preceded by an instruction (prompt) to GPT-4o. In this prompt, GPT-4o is asked to determine whether the output contains hallucinations, and we expect a single-word response of either “Yes” or “No.” An example of such an instruction for the question answering task is shown in Table 6.
For human annotation, we asked three team members with at least upper-intermediate English proficiency to independently annotate approximately 100 samples from each dataset. We selected samples where all annotators reached a consensus and considered these annotations the ground truth hallucination labels.
To further evaluate GPT-4o, we conducted automatic annotation using several variations of prompts, each reformulating the task for GPT-4o, including zero-shot and few-shot versions. We then compared these annotations to the actual hallucination labels. The results, presented in Table 7, demonstrate a consistent alignment between GPT-4o’s annotations and those made by humans, regardless of the specific prompt. This consistency confirms the robustness of our approach to the exact form of instruction.
Based on these findings, we prefer automated annotation as a cost-effective and efficient alternative to human experts.
Annotation: general pipeline. CoQA and SQuAD contain questions paired with ground-truth answers. To minimize false positives in labeling, we employed a two-step verification process:
  1. 1.
    Rouge-L scoring: we computed Rouge-L scores (using the evaluate library) between the model’s response and the ground-truth answers.
  2. 2.
    Substring matching: we checked whether any ground-truth answer was a substring of the response.
Responses with a Rouge-L score of 1 (exact match) were labeled as grounded. Those meeting both of the following criteria were flagged as potential hallucinations:
  • Rouge-L score $\leq 0.3$ (following  Kuhn et al. (2023));
  • no ground-truth answer appears as a substring.
These candidate hallucinations were then reviewed by GPT-4o, and only confirmed cases were finally labeled as hallucinations.
For XSum, where reference summaries are more complex than the ground truth answers in SQuAD/CoQA, we bypassed Rouge-L filtering and relied solely on GPT-4o for annotation.
Detailed statistics for each dataset can be seen in Table 8. The number of samples in the datasets varies across models, as we tried to maintain a balance of hallucinated and grounded responses, ensure sample cleanness, and minimize mislabeling. The procedure outlined above selects a different number of objects in a sample depending on the quality of the model’s responses.
Model CoQA SQuAD XSum
Hal. Grounded Hal. Grounded Hal. Grounded
Mistral-7B 776 776 311 389 301 448
LLaMA-2-7B 375 375 357 235 239 507
LLaMA-2-13B 279 384 314 436 208 522
LLaMA-3.1-8B 189 200 350 400 243 407
Qwen2.5-7B 124 183 215 249 194 556
Table 8: Datasets statistics. Number of hallucinated and grounded samples of each model.
Appendix D Other experiment results
D.1 Alternative attention map-based features for hallucination detection
Figure 8: $\Delta_{ij}$ values for $ij$-th heads, MS MARCO vs CoQA. Vertical axis corresponds to the difference on dataset (B), horizontal — to the one on dataset (A). The heads that segregate samples best are highlighted in pink. Model names for a row are on the left side.
Table 9: ROC-AUC values of supervised classifiers on top of various set of features. TOP-1 results are highlighted with bold font, while TOP-2 are underlined.
Features MS MARCO CoQA
Mistral-7B
Standard topological 0.67 0.69
Sparsity ratio 0.66 0.7
Entropy 0.75 0.77
Wasserstein 0.77 0.73
Spectral norm 0.73 0.72
Attention to <s> 0.65 0.61
MTop-Div 0.86 0.98
LLaMA-2-7B
Naive topological 0.69 0.7
Sparsity ratio 0.49 0.61
Entropy 0.38 0.68
Wasserstein 0.73 0.6
Spectral norm 0.49 0.64
Attention to <s> 0.62 0.64
MTop-Div 0.75 0.96
In our preliminary experiments for developing an attention maps-based hallucination detector, we trained classifiers using topological features previously applied to other NLP tasks Kushnareva et al. (2021); Cherniavskii et al. (2022), as well as traditional attention map characteristics. As standard topological features, we considered barcode-based features, such as the sum of bar lengths in persistence diagrams, and naive topological features, including the average vertex degree in attention graphs. For traditional attention-based features, we used sparsity ratio, attention entropy, and spectral norm Kobayashi et al. (2020); Vig and Belinkov (2019); Ji et al. (2021). We also considered Wassersein distances between the persistent diagrams Edelsbrunner and Harer (2010) of a context and a response subgraphs as an alternative way to assess their similarity. Finally, we analyzed the average attention to <s> token as we discovered that hallucination-aware heads often attend to it when the hallucination is present (see 4).
To determine the most informative features for hallucination detection, we trained supervised classifiers on concatenated features from all layers and heads and compared them to classifiers trained only on MTop-Div values under the same conditions. The results are presented in Table 9.
While the classifier based on MTop-Div values significantly outperforms alternative approaches, computing these values across all layers and attention heads is highly computationally expensive. To address this, we developed TOHA — a more efficient alternative by aggregating MTop-Div values from only a subset of the “hallucination-aware” attention heads in the model.
D.2 Other metrics for the head selection procedure
We also investigated alternative attention-map-based scores — including entropy, spectral norm, and the Wasserstein distance between the persistent diagrams of prompts and responses — for selecting specialized attention heads.Following the same pipeline as in TOHA, we computed the average distances between hallucinated and grounded samples using alternative proximity metrics. The results, presented in Figure 8, reveal that the most segregating heads for MS MARCO do not generalize to the CoQA dataset. This suggests that our proposed MTop-Div metric is better suited for the task compared to existing solutions.
Appendix E Implementation details
In this section, we describe the key implementation choices.
  • For EigenScore, we used the last token representation to embed sentences, as suggested in Chen et al. (2024). We took outputs from the 16th layer, since middle layers were shown to contain the most factual information Sky et al. (2024); Azaria and Mitchell (2023).
  • For methods that rely on multiple generations, we generated 20 samples per input, following recommendations from Manakul et al. (2024); Chen et al. (2024).
  • For SelfCheck-GPT, we used its NLI-based variant.
  • For LLM-Check, we considered its white-box attention score modification, as it works in a setting similar to ours.
  • For Haloscope, we reserved 150 samples for hyperparameter tuning and treated the rest as unlabeled data. For the CoQA dataset, we added several thousand extra generations during training, as in the original paper. Other datasets were used as-is, since their sizes are comparable with those in Haloscope’s experiments on the TruthfulQA dataset.
  • For TOHA’s head selection, we similarly used 150 annotated samples. The topological divergences were calculated using ripser++ library Zhang et al. (2020), MIT license.
In our experiments, we used 60/15/25 train/val/test split. All the obtained results were averaged over five runs. All experiments were carried out using NVidia L40.
Appendix F Use of scientific artifacts & AI assistants
CoQA contains passages from seven domains under the following licenses: Literature and Wikipedia passages are shared under CC BY-SA 4.0 license; Children’s stories are collected from MCTest which comes with MSR-LA license; Middle/High school exam passages are collected from RACE which comes with its own license; News passages are collected from the DeepMind CNN dataset which comes with Apache license. SQuAD dataset comes under CC BY-SA 4.0 license. RAGTruth dataset comes under MIT license. XSum dataset comes under MIT license.
We used all the artifacts as it was intended by the corresponding licenses. No personal information or offensive content is contained in the considered datasets.
The original text of this paper was spell- and grammar-checked and slightly smoothed out using Grammarly.
Appendix G Potential risks
  1. 1.
    Ethical risks from deployment: overconfidence in TOHA’s scores could lead to unchecked LLM outputs in high-stakes scenarios (e.g., healthcare). TOHA should be frame as a "warning system" rather than a definitive filter, and advocate for human review.
  2. 2.
    Attention manipulation attacks: adversarial prompts could artificially alter attention patterns, evading detection.

### v3

@@
 Hallucination, i.e., generating factually incorrect content, remains a critical challenge for large language models (LLMs).
-We introduce TOHA [^1] The code of the proposed method and the considered baselines is available at https://anonymous.4open.science/r/tda4hallu-BED5, a TOpology-based HAllucination detector in the RAG setting, which leverages a topological divergence metric to quantify the structural properties of graphs induced by attention matrices.
+We introduce TOHA [^1] The code of the proposed method and the considered baselines is available at https://anonymous.4open.science/r/tda4hallucinations-C449, a TOpology-based HAllucination detector designed for RAG scenarios, which leverages a topological divergence metric to quantify the structural properties of graphs induced by attention maps.
 Examining the topological divergence between prompt and response subgraphs reveals consistent patterns: higher divergence values in specific attention heads correlate with hallucinated outputs, independent of the dataset.
-Extensive experiments — including evaluation on question answering and summarization tasks — show that our approach achieves state-of-the-art or competitive results on several benchmarks while requiring minimal annotated data and computational resources.
+Extensive experiments, including evaluation on question answering and summarization tasks, show that our approach achieves state-of-the-art or competitive results on several benchmarks while requiring minimal annotated data and computational resources.
 Our findings suggest that analyzing the topological structure of attention matrices can serve as an efficient and robust indicator of factual reliability in LLMs.
@@
 Hallucination Detection in LLMs with Topological Divergence on Attention Graphs
-Alexandra Bazarova1, Aleksandr Yugay1, Andrey Shulga1, Alina Ermilova1, Andrei Volodichev1, Konstantin Polev2, Julia Belikova2, Rauf Parchiev 2, Dmitry Simakov2, Maxim Savchenko2, Andrey Savchenko2, Serguei Barannikov1, 3, Alexey Zaytsev1, 1Skolkovo Institute of Science and Technology, 2Sber AI Lab, 3CNRS, Universite Paris Cite, Correspondence: a.bazarova@skoltech.ru
+Alexandra Bazarova1, Aleksandr Yugay1, Andrey Shulga1, Alina Ermilova1, Andrei Volodichev1, Konstantin Polev2, Julia Belikova2, Rauf Parchiev 2, Dmitry Simakov2, Maxim Savchenko2, Andrey Savchenko2, Serguei Barannikov1, 3, Alexey Zaytsev1, 1Applied AI Institute, 2SB AI Lab, 3CNRS, Universite Paris Cite, Correspondence: bazarovaai.239@gmail.com
 **footnotetext: Equal contribution.
@@
 (2024).
-However, these models are prone to generate so-called hallucinations, i.e., content that is factually or contextually incorrect Huang et al.
+To ensure factual reliability, modern LLMs are often combined with retrieval-augmented generation (RAG) technique, which integrates relevant external knowledge from diverse databases directly into the generation process Lewis et al.
 (2020).
-Detecting hallucinations is crucial for safely deploying LLMs in sensitive fields since erroneous outputs may seriously harm user trust.
+Despite these improvements, LLMs remain prone to producing so-called hallucinations, i.e., content that is factually or contextually incorrect Huang et al.
+(2023); Li et al.
+(2024).
+Detecting hallucinations is crucial for safe deployment of LLMs in sensitive fields since erroneous outputs may seriously harm user trust.
 An effective detector would therefore expand the scope of LLM applications while mitigating risks Gao et al.
 (2024).
-Multiple methods address this problem Huang et al.
-(2023); Sahoo et al.
-(2024), though many face significant practical constraints.
-A common limitation is the reliance on large annotated datasets Sky et al.
-(2024); Azaria and Mitchell (2023); Chuang et al.
-(2024), which are rarely available publicly Zhang et al.
-(2023) and require extensive annotation effort for each new model released.
-Another popular approach depends on generating multiple samples for scoring Manakul et al.
-(2024); Chen et al.
-(2024); Farquhar et al.
-(2024), which increases computational costs substantially.
-Figure 1: Difference between average topological divergence values for hallucinated and grounded samples per attention head/layer, evaluated on MS MARCO and CoQA datasets.
-A lighter color corresponds to a greater difference.
-Green frames highlight the heads that separate samples best.
-The same attention heads assign greater divergence values to the hallucinated samples in both datasets.
-Model: Mistral-7B-Instruct-v0.1.
-We address these challenges by introducing TOHA (TOpology-based HAllucination detector), a training-free method for the retrieval-augmented generation (RAG) setting Gao et al.
+Multiple methods have been proposed to tackle this problem Sahoo et al.
+(2024); Shorinwa et al.
 (2025).
-Following the prior work Du et al.
-(2024), TOHA requires minimal annotated data (just $50$ annotated samples suffice for reliable detection, see Figure 4) while avoiding the computational overhead of multiple generations, making it both data- and compute-efficient.
-The core insight behind TOHA is that hallucinated answers tend to have weaker connections to the context than grounded ones in the RAG setting.
-TOHA formalizes this by analyzing attention graphs — complete graphs induced by LLM attention maps, a representation previously used in other NLP tasks Kushnareva et al.
+However, they often face significant practical constraints, such as the scarcity of annotated datasets Zhang et al.
+(2023) required for supervised methods Sky et al.
+(2024); Orgad et al.
+(2025), the high computational cost of generating multiple additional samples Chen et al.
+(2024); Hou et al.
+(2025), or the inability of LLMs’ output probabilities to fully represent the model’s true uncertainty Fadeeva et al.
+(2024); Shelmanov et al.
+(2025).
+These challenges can be addressed by utilizing the inner states of LLMs, which are informative for the hallucination detection problem Azaria and Mitchell (2023); Sriramanan et al.
+(2024); Gekhman et al.
+(2025).
+We introduce TOHA (a TOpology-based HAllucination detector), a training-free method designed for RAG scenario that leverages the structure of LLM attention maps to identify hallucinations.
+Our method requires minimal annotated data while avoiding the computational overhead of multiple generations, which makes TOHA both data- and compute-efficient.
+The core insight behind TOHA is that the structure of prompt-response interconnections within an LLM’s attention mechanism should be indicative of hallucination presence.
+TOHA formalizes this idea by analyzing attention graphs — complete weighted graphs derived from LLM attention maps, the representation prior used for topological data analysis (TDA) in NLP Kushnareva et al.
 (2021); Tulchinskii et al.
 (2023).
-Unlike prior works that rely on simplistic classifiers over basic graph properties Proskurina et al.
-(2023a); Cherniavskii et al.
-(2022), TOHA advances the approach by computing the topological dissimilarity between the attention subgraphs of the model’s response and the given context.
-This dissimilarity, measured via our adaptation of Manifold Topology Divergence  Barannikov et al.
-(2021) for the graph setting, quantifies the “strength” of context-response ties: higher values indicate weaker links and thus likely hallucinations.
-We prove key stability properties for this metric, ensuring its reliability as a hallucination score.
+Unlike the existing attention-based methods, which either treat all attention heads as equally important Sriramanan et al.
+(2024); Binkowski et al.
+(2025) or ignore the geometric structure of attention maps Sun et al.
+(2025); Vazhentsev et al.
+(2025), our method leverages a small set of attention heads selected by a topological criterion designed to identify hallucinations.
+Specifically, TOHA is based on the $\operatorname{MTop-Div}_{G}(R,P)$, our adaptation of Manifold Topology Divergence  Barannikov et al.
+(2021) for the graph setting, which quantifies the dissimilarity between the prompt ($P$) and response ($R$) token sets in the attention graph.
+We demonstrate that this score not only reflects the geometry but also measures informational novelty of the response in relation to the prompt, thus being well-suited for hallucination detection in RAG scenarios.
 Figure 1: a) An attention map.
@@
 To keep figure neat, we only plot the edges with an attention score of no less than $0.15$.
-c) The minimum spanning forest attaching $R$ to $P$.
-Through analysis of divergence patterns across different heads, we identified a subset of attention heads that consistently assign higher divergence scores to hallucinated samples (see Figure 1), revealing their implicit “awareness” of hallucinations.
-TOHA utilizes the average divergence values from these specific heads as hallucination scores.
-Crucially, these heads exhibit consistent behavior across different datasets, enabling strong transferability of our method.
-Our main contributions can be summarized the following:
-  • We propose TOHA, a training-free method based on the topological divergences of attention graphs.
-While efficient — TOHA operates up to an order of magnitude faster than methods of comparable quality and requires minimal annotated data — our method demonstrates strong in-domain performance and maintains domain transferability across different tasks.
-  • The existence of hallucination-aware attention heads is discovered: calculating topological divergences from just six specific heads is enough for reliable hallucination detection, irrespective of the dataset.
-  • Our experiments show TOHA consistently matches or exceeds state-of-the-art performance on all benchmarks when applied to modern open-source LLMs of varying scales (7B to 13B parameters).
+c) The minimum spanning forest attaching $R$ to $P$ and the corresponding $\operatorname{MTop-Div}$ value.
+Through analysis of divergence values across attention heads, we identified a subset of heads that consistently yield greater scores for hallucinated samples (see Figure 2), irrespective of the dataset.
+TOHA’s final hallucination score is the average $\operatorname{MTop-Div}_{G}(R,P)$ value from these “hallucination-aware” heads.
+We discover that some of these heads are associated with copying behavior, which is aligned with prior findings Sun et al.
+(2025).
+Our main contributions can be summarized as follows:
+  • We introduce TOHA, a training-free framework for hallucination detection that leverages the topology of LLM attention graphs.
+Our method operates up to an order of magnitude faster than methods of comparable quality and requires minimal annotated data.
+  • Central to TOHA is $\operatorname{MTop-Div}_{G}(R,P)$, an adaptation of Manifold Topology Divergence for graphs, which quantifies the topological dissimilarity between prompt and response sets of tokens.
+We demonstrate that this score characterizes the novelty of the response with respect to the prompt, making it highly effective for detecting hallucinations in RAG systems.
+  • By analyzing the proposed score, we discover the existence of “hallucination-aware” attention heads, which consistently yield greater divergence values for hallucinated samples across different datasets.
+This finding ensures TOHA’s efficiency and solid cross-domain transferability: averaging $\operatorname{MTop-Div}_{G}(R,P)$ for just ten heads is enough for robust hallucination detection.
+  • Our experiments show that TOHA consistently matches or exceeds state-of-the-art performance across multiple benchmarks when applied to open-source LLMs of varying scales, including 7B and 13B parameters.
 2 Background
-2.1 Attention matrix as a weighted graph
+2.1 Attention Map as a Weighted Graph
 Modern LLMs are mainly based on the self-attention mechanism, introduced in Vaswani et al.
@@
 $\operatorname{Attention}(Q,K,V)=W(Q,K)V,$ (1)
-where $W(Q,K)$ is an attention matrix
-$\operatorname{W}=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right),$ (2)
-and each entry $w_{ij}$ in it captures how strongly token $i$ attends to token $j$, $i\geq j$ for a decoder, with larger $w_{ij}$ indicating closer relationship.
-An attention matrix $W$ can be reframed as a complete weighted graph $G$ where tokens are vertices and weights $w_{ij}$ represent the strength of connections between them.
-From the perspective of topological data analysis, however, it is more convenient to consider these weights as pseudo-distances rather than correlation measures.
-Hence, we reassign the edge weights of such a graph to equal $1-w_{ij}$, creating what we call attention graphs.
-In the generation process, the vertices of such graphs naturally partition into two distinct subsets: prompt tokens $P$ and response tokens $R$ generated by the model (see Figure 2b).
-This split allows us to analyze the topological relationships between input and output content.
+where $W(Q,K)$ is an attention map
+$\operatorname{W}(Q,K)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right),$ (2)
+and each entry $w_{ij}=W_{ij}(Q,K)$ captures how strongly token $i$ attends to token $j$, with greater values indicating closer relationship.
+We consider decoder-only LLMs, for which the attention maps are lower-triangular.
+An attention map can be reframed as a complete undirected weight graph $G$ with edge weights $1-w_{ij}$ to represent pseudo-distances between tokens.
+We call $G$ an attention graph.
+It naturally partitions into prompt ($P$) and response ($R$) tokens (see Figure 1b).
+We analyze the topological relationships between these node subsets, which, as we assume, should be indicative of hallucinations in the RAG scenario.
 2.2 Manifold Topology Divergence
-One way to compare two data manifolds, $\mathcal{M}$ and $\mathcal{N}$, approximated by point clouds $M$ and $N$ is the $\operatorname{MTop-Div}(M,N)$ topological measure Barannikov et al.
+The Manifold Topology Divergence, or $\operatorname{MTop-Div}$, was proposed in Barannikov et al.
 (2021).
-This divergence is based on the $\operatorname{Cross-Barcode}(M,N)$, which is a set of intervals $\{(b_{i},d_{i})\}_{i=1}^{n}$ corresponding to “births” and “deaths” of independent topological features that distinguish the point cloud $N$ from the union $M\cup N$.
-The further $\operatorname{Cross-Barcode}(M,N)$ is from an empty set, the more the data manifold $\mathcal{N}$ differs from $\mathcal{M}$ in its topological structure.
-To measure the distance from an empty set, the sum of lengths of intervals in $\operatorname{Cross-Barcode}(M,N)$ was taken in loc.
-cit.:
-$\operatorname{MTop-Div}(M,N)=\sum\limits_{i=1}^{n}|d_{i}-b_{i}|.$
+$\operatorname{MTop-Div}(M,N)$ quantifies the difference between two data manifolds $\mathcal{M}$ and $\mathcal{N}$, approximated by point clouds $M$ and $N$.
+It is computed as the sum of interval lengths in the $\operatorname{Cross-Barcode}(M,N)$, a set of intervals representing topological features distinguishing $N$ from $M\cup N$.
+A larger divergence indicates a greater topological difference between the manifolds.
+For more details on topological data analysis, see Appendix C.
 3 Method
-Given an attention matrix for the (prompt + response) text, we construct the attention graph, imitating a data manifold of the text, and study its relation with the weighted subgraph, imitating the data submanifold of the prompt.
-We measure the topological divergence between these graphs, assuming responses that are consistent with the context would result in lower divergence values.
-3.1 $\operatorname{MTop-Div}$ for attention graphs
-The $\operatorname{MTop-Divergence}(M,N)$ measure was originally developed for data manifolds, relying on metric space properties to some extent.
-We adapt this concept to quantify divergence between attention graphs and their subgraphs, where traditional metric axioms do not apply, while preserving the measure’s comparative utility.
-Let $R$ and $P$ be the sets of response and the prompt vertices in the attention graph $G$.
-We set to zero the edge weights between the $P$ vertices, denote $w_{(R\cup P)/P}$ the resulting matrix of edge weights, and define $\operatorname{Cross-Barcode}_{i}(R,P)$ as the $i$-th homology barcode of the Vietoris-Rips simplicial complex $VR_{\alpha}(G,w_{(R\cup P)/P})$.
-We define $\operatorname{MTop-Div}(R,P)$ as the total sum of interval lengths in $\operatorname{Cross-Barcode}_{0}(R,P)$, where we consider the 00-dimensional homology group $H_{0}$.
-In this context, we can prove that this score is equivalent to the total edge length of the minimum spanning forest (MSF) connecting $R$ to $P$.
-Properties of MTop-Div for attention graphs.
-Here, we only list the properties relevant to hallucination detection; for more properties and proofs, see Appendix A.
+This section introduces the $\operatorname{MTop-Div}_{G}(R,P)$ score, which is designed to quantify the topological divergence between the prompt and response subgraphs in attention maps.
+We demonstrate that not only can it be interpreted as a geometric characteristic of the attention graph, but also as an information-theoretic measure of the novelty of the response in relation to the prompt.
+With $\operatorname{MTop-Div}_{G}(R,P)$, we identify “hallucination-aware” heads that consistently separate hallucinated samples from grounded, irrespective of dataset.
+This finding allows us to formulate the TOHA algorithm (Algorithm 1), which computes a final hallucination score by averaging the divergence values from these specific heads.
+3.1 $\operatorname{MTop-Div}$ for Attention Graphs: Definition
+Let $R$ and $P$ be the response and prompt vertex sets in an attention graph $G$.
+After zeroing edge weights between the $P$ nodes, we compute the 0-th order homology barcode $\mathcal{B}_{0}$ of the Vietoris-Rips simplicial complex of the modified graph.
+Essentially, this barcode tracks the evolution of connected components as we threshold the graph edges (see Appendix C for details).
+Our proposed topological divergence is a sum of lengths of $\mathcal{B}_{0}$ intervals:
+$\operatorname{MTop-Div}_{G}(R,P)=\sum\limits_{\scriptstyle[b_{i},d_{i}]\in\mathcal{B}_{0}}|d_{i}-b_{i}|.$
+This score can be interpreted from two perspectives: geometric (as the length of a minimal spanning forest in the attention graph) and information-theoretic (as a measure of the response novelty in the space induced by query-key matrices).
+We explore these two interpretations in the following subsections.
+3.2 $\operatorname{MTop-Div}_{G}(R,P)$ and Geometry of Attention Graph
+Formally, we can prove the following property (see proof in Appendix C):
 Proposition 3.1.
-The following holds for any attention graph $G$ with vertex set $V_{G}$ and its complementary vertex subsets $P,R$, where $P\cup R=V_{G}$ and $P\cap R=\varnothing$.
-1. (Formula.) $\operatorname{MTop-Div}(R,P)$ value equals the length of the MSF attaching $R$ to $P$.
-2. (Stability.) If the weights of $G$ change by no more than $\varepsilon$, then the corresponding $\operatorname{MTop-Div}(R,P)$ changes by no more than $\delta=\varepsilon|R|$.
-3.
-(Connection with hallucinations.) The normalized divergence value $\frac{1}{|R|}\operatorname{MTop-Div}(R,P)=0$ iff the MSF attaches every response token to a prompt token by a subtree with attention weights $=1$.
-The stability property guarantees that similar attention patterns yield similar hallucination scores, making the metric’s behavior consistent and predictable.
-The latter captures the intuitive relationship between divergence and response quality: well-grounded responses (closely tied to the context) produce small divergence values, while hallucinations (occurring when evidence is missing) lead to smaller attention weights and consequently larger divergence values.
-Together, these properties enable the metric to reliably measure how strongly a response connects to its context while remaining robust to minor attention variations.
+Consider an attention graph $G$ with vertex set $V_{G}$ and its complementary vertex subsets $P,R$, where $P\cup R=V_{G}$ and $P\cap R=\varnothing$.
+$\operatorname{MTop-Div}(R,P)$ value equals the length of the minimal spanning forest (MSF) attaching $R$ to $P$.
+This proposition’s geometric meaning is illustrated in Figure 1.
+While the connections between prompt nodes (Figure 1b) are semantically and syntactically meaningful, we hypothesize they primarily introduce noise for hallucination detection (see Section 4.4 for the corresponding experiment).
+After setting these distances to zero, we construct an MSF on the modified graph (Figure 1c).
+The sum of the lengths of the remaining edges in this MSF is precisely the value of $\operatorname{MTop-Div}_{G}(R,P)$.
+3.3 $\operatorname{MTop-Div}_{G}(R,P)$ and Perplexity in the Query-Key Space
+Consider an attention graph $G$.
+It can be interpreted as a non-metric***The $Q^{T}K$ scalar product from self-attention cannot be formally interpreted as a metric: for one, it does not adhere to the triangle’s inequality.
+space, with “distances” induced by $Q$ and $K$ matrices from the attention mechanism.
+In order to quantify some notion of entropy in this non-metric space, the most intuitive way is to measure the MST length Müller et al.
+(2012) as a proxy for token dispersion.
+More precisely, the entropy estimate given the MST of length $L$ on points $\mathcal{X}$ is the following:
+$H_{MST}(\mathcal{X})=d\log L-(d-1)\log n+\log\beta_{d},$
+where $|\mathcal{X}|=n$, $d$ is the intrinsic dimensionality of the data, and $\beta_{d}$ is some data-independent constant.
+We can show that
+$\operatorname{MTop-Div}_{G}(R,P)\geq L_{MST}(R\cup P)-L_{MST}(P),$
+where the first term corresponds to the MST length of the complete attention graph $G$, and the second to the length of the MST spanning the prompt tokens $P$.
+Therefore, our score has the following lower bound:
+$\operatorname{MTop-Div}_{G}(R,P)\geq\alpha 2^{H(P\cup R)}-\beta 2^{H(P)},$
+$\alpha,\beta>0$, or, in terms of the perplexity function,
+$\operatorname{MTop-Div}_{G}(R,P)\geq\alpha\mathrm{PPL(P\cup R)}-\beta\mathrm{PPL(R)}.$
+Perplexity measures how “surprising” a piece of data is.
+Therefore, the difference between the perplexity values quantifies the novelty of the response with respect to the prompt: greater values suggest that new information is present, which, in the RAG scenario, is considered to be a hallucination.
+Thus, our $\operatorname{MTop-Div}_{G}(R,P)$ should be an effective statistic for identifying hallucinated responses.
 3.4 Hallucination-Aware Heads
-We hypothesize, inspired by prior investigations in LLM interpretability Voita et al.
-(2019); Gould et al.
-(2024), that particular attention heads exhibit distinct patterns related to hallucinations.
-To identify such heads, we analyzed head-specific topological divergences as follows.
 Denote by $h_{ij}$ the $j$-th attention head from the layer $i$.
@@
 We examined typical values of the average distance between hallucinated and grounded training examples for different heads and layers:
-$\Delta_{ij}=\frac{1}{|S_{\mathrm{hallu}}|}\sum\limits_{s\in S_{\mathrm{hallu}}}d_{ij}(s)-\frac{1}{|S_{\mathrm{gr}}|}\sum\limits_{s\in S_{\mathrm{gr}}}d_{ij}(s),$
+$\Delta_{ij}=\frac{1}{|S_{\mathrm{hallu}}|}\sum\limits_{s\in S_{\mathrm{hallu}}}d_{ij}(s)-\frac{1}{|S_{\mathrm{gr}}|}\sum\limits_{s\in S_{\mathrm{gr}}}d_{ij}(s),$ (3)
 where $S_{\mathrm{hallu}}$ stands for all hallucinated samples from the training set, $S_{\mathrm{gr}}$ stands for all grounded training samples, and
-$d_{ij}(s)=\frac{1}{|R_{ij}^{s}|}\operatorname{MTop-Div}(R^{s}_{ij},P^{s}_{ij}).$
-The sample differences obtained for three datasets are displayed in Figure 3.
-Each dot represents an individual attention head, with its $x$-coordinate indicating its $\Delta_{ij}$ value on dataset (A) and its $y$-coordinate the corresponding value on dataset (B).
-For each dataset pair, we highlight the top three most separating heads for dataset (A) in pink.
-Notably, these heads consistently appear in the upper-right corner of the plot, indicating that they also exhibit strong separation on dataset (B).
-This observation suggests that these heads are inherently attuned to hallucination patterns, regardless of the dataset.
-Figure 3: $\Delta_{ij}$ values for $ij$-th heads.
+$d_{ij}(s)=\frac{1}{|R_{ij}^{s}|}\operatorname{MTop-Div}_{G^{s}_{ij}}(R^{s}_{ij},P^{s}_{ij}).$
+Figure 2 displays sample differences $\Delta_{ij}$ across three datasets, with each marker representing some attention head.
+We discovered that the same four (for Mistral-7B) and three (for Llama-2-7B) heads, highlighted in pink, demonstrate similar behaviour across the datasets: they consistently appear in the upper-right corner, indicating strong separation between hallucinated and grounded samples, irrespective of the dataset.
+This finding suggests that there are specific attention heads somehow “aware” of the presence of hallucination, which is captured by the proposed $\operatorname{MTop-Div}_{G}(R,P)$.
+Figure 2: $\Delta_{ij}$ values for ijij-th heads.
 Vertical axis corresponds to the difference on dataset (B), horizontal — to the one on dataset (A).
-The heads that separate samples best are highlighted in pink.
-Model names for a row are on the left side, datasets: MS MARCO, CNN/DM + Recent News, CoQA.
+The heads that separate samples best are highlighted in pink; their (layer, head) positions are reflected in the legend.
 3.5 TOHA
-The existence of universal hallucination patterns in the attention heads underlies our efficient method TOHA, detailed in Algorithm 1.
-It uses two small, annotated probe sets, $S_{h}$ (containing hallucinated samples) and $S_{g}$ (containing grounded samples), to rank model heads by their separation capability based on their $\Delta_{ij}$ values, where $ij$ denotes a head index, and select the most relevant ones.
-In our experiments, the combined size of the probe sets is kept small (see Figure 4 for the number of the required samples analysis).
-During testing, hallucination scores are computed as the average topological divergence from the top $N_{\mathrm{opt}}$ heads, where $N_{\mathrm{opt}}$ is a hyperparameter tuned on the set $V=S_{h}\cup S_{g}$.
-For computational efficiency, we limit $N_{\mathrm{opt}}$ to a maximum of 6 in all experiments.
+The existence of “hallucination-aware” attention heads underlies our method, which is detailed in Algorithm 1.
+TOHA uses two annotated probe sets ($S_{h}$ for hallucinations, $S_{g}$ for grounded samples) to rank attention heads by their separation capability $\Delta_{ij}$ and select the top $N_{\mathrm{opt}}$ ones with the combined probe size $|S_{h}\cup S_{g}|$ kept small (see Figure 3).
+During testing, the final hallucination score for test samples $T$ equals the average topological divergence from the selected $N_{\mathrm{opt}}$ heads.
+For computational efficiency, we limit $N_{\mathrm{opt}}$ to a maximum of $N_{\max}=10$ in all experiments.
 Algorithm 1 TOHA Algorithm
-$d_{ij}(s)$ — topological divergences for samples;
-$S_{h},\,S_{g}$ — probe sets;
-$V=S_{h}\cup S_{g}$ — validation set;
-$T$ — test set;
-$N_{\mathrm{max}}$ — max number of selected heads.
-procedure TOHA heads selection
-     $\Delta_{ij}\leftarrow\frac{1}{|S_{h}|}\sum\limits_{s\in S_{h}}d_{ij}(s)-\frac{1}{|S_{g}|}\sum\limits_{s\in S_{g}}d_{ij}(s)$
-     $H\leftarrow\mathrm{sort\_descending}(h_{ij},\text{key}=\Delta_{ij})$
-     $N,N_{\mathrm{opt}}\leftarrow 1,1$
-     $H_{\mathrm{subset}}\leftarrow\varnothing$triangleright Optimal heads set.
-     $\text{AUROC}_{\max}\leftarrow 0$
-     $p_{s}=0,s\in V$triangleright Hallucination scores.
-     while $N\leq N_{\max}$ do
-         $H_{\mathrm{subset}}\leftarrow H_{\mathrm{subset}}\cup\{h_{N}\}$
-         for $s\in V$ do
-              $p_{s}\leftarrow\frac{N-1}{N}p_{s}+\frac{1}{N}d_{h_{N}}(s)$
-         end for
-         $\text{auroc}\leftarrow\text{AUROC}(\{y_{s}\}_{s\in V},\{p_{s}\}_{s\in V})$
-         if $\text{auroc}>\text{AUROC}_{\max}$ then
-              $\text{AUROC}_{\max}\leftarrow\text{auroc}$
-              $N_{\mathrm{opt}}\leftarrow N$
-         end if
-         $N\leftarrow N+1$
-     end while
-end procedure
-procedure TOHA prediction
-     for $s\in T$ do $\triangleright$ Prediction on the test set.
-         $p_{s}\leftarrow\frac{1}{N_{\mathrm{opt}}}\sum\limits_{i=1}^{N_{\mathrm{opt}}}d_{h_{i}}(s)$
-     end for
-end procedure
-4 Experiments
-Datasets.
-We evaluated our approach on four datasets: RAGTruth Niu et al.
-(2023) (we considered its’ two separate benchmarks: QA based on MS MARCO Nguyen et al.
-(2016) and summarization based on CNN/DM Nallapati et al.
-(2016)) combined with news articles from an unnamed news platform), CoQA Reddy et al.
-(2019), SQuAD Rajpurkar et al.
-(2016), and XSum Narayan et al.
+[…]
+2:Hallucination scores $\{p_{s}\}$ for $s\in T$
+3:
+4:function HeadsSelection
+5:  for each head $h_{ij}$ do
+6:   Calculate $\Delta_{ij}$ according to eq. (3)
+7:  end for
+8:  $H\leftarrow$ sort all heads by $\Delta_{ij}$ (descending)
+9:  $N_{\text{opt}},\text{AUROC}_{\max}\leftarrow 1,0$
+10:  Initialize $p_{s}\leftarrow 0$ for all $s\in V$
+11:  for $N=1$ to $N_{\max}$ do
+12:   for each $s\in V$ do
+13:     $p_{s}\leftarrow\frac{N-1}{N}p_{s}+\frac{1}{N}d_{h_{N}}(s)$
+14:   end for
+15:   $\text{auroc}\leftarrow\text{AUROC}(\{y_{s}\}_{s\in V},\{p_{s}\}_{s\in V})$
+16:   if $\text{auroc}>\text{AUROC}_{\max}$ then
+17:     $\text{AUROC}_{\max}\leftarrow\text{auroc};\ N_{\text{opt}}\leftarrow N$
+18:   end if
+19:  end for
+20:  return $N_{\text{opt}}$
+21:end function
+22:
+23:function Prediction($N_{\text{opt}}$)
+24:  for each $s\in T$ do
+25:   $p_{s}\leftarrow\frac{1}{N_{\text{opt}}}\sum_{i=1}^{N_{\text{opt}}}d_{h_{i}}(s)$
+26:  end for
+27:  return $\{p_{s}\}$
+28:end function
+4 Results
+4.1 Experiment Setting
+Datasets and models.
+We consider five tasks that evaluate question answering and summarization abilities of LLMs: two benchmarks of RAGTruth Niu et al.
+(2023) — long-form QA dataset MS MARCO and summarization dataset CNN/DM; conversational QA dataset CoQA Reddy et al.
+(2019), reading comprehension dataset SQuAD Rajpurkar et al.
+(2016), and extreme summarization dataset XSum Narayan et al.
 (2018).
-The RAGTruth dataset consists of manually annotated responses of several LLMs in the RAG setting.
-The annotations are word-level; we, in turn, predict response-level labels, considering a response hallucinated if it contains at least one hallucination span.
-For the three latter datasets, we sampled LLM responses and annotated them automatically with GPT-4o Hurst et al.
-(2024).
-Consistency with Human-GPT-4o label (Table 7) validated this approach, matching prior work Bavaresco et al.
-(2024); see Appendix C for more details.
-Models.
-We used five popular open-source LLMs: LLaMA-2-7B-chat, LLaMA-2-13B-chat, LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct-v0.1, and Qwen2.5-7B-Instruct.
-Note that the RAGTruth dataset does not contain responses for LLaMA-3.1-8B and Qwen-2.5-7B; therefore, we only conducted experiments on SQuAD, CoQA, and XSum for these models.
+For more details, see Appendix E.
+We employ five popularly adopted open-source LLMs with accessible inner states: LLaMA-2-7B-chat, LLaMA-2-13B-chat, LLaMA-3.1-8B-Instruct, Mistral-7B-Instruct-v0.1, and Qwen2.5-7B-Instruct.
+As the RAGTruth dataset does not contain responses for LLaMA-3.1-8B and Qwen-2.5-7B, we conducted experiments on SQuAD, CoQA, and XSum for these models.
 Baselines.
-We compare TOHA with seven baselines: perplexity Ren et al.
-(2023), max entropy Fadeeva et al.
+We compare TOHA with a comprehensive set of eight baselines: uncertainty-based perplexity Ren et al.
+(2023) and max entropy Fadeeva et al.
+(2024); inner states-based ReDEEP Sun et al.
 (2025), HaloScope Du et al.
-(2024), LLM-Check Sriramanan et al.
-(2024), semantic entropy Farquhar et al.
+(2024), and LLM-Check Sriramanan et al.
+(2024); consistency-based semantic entropy Farquhar et al.
 (2024), EigenScore Chen et al.
@@
 (2024).
-Appendix E provides information on implementation details.
-Table 1: ROC AUC ()($\uparrow$)( ↑ ) of hallucination detection techniques for three LLMs. The best results for each model are highlighted in bold, and the second best are underlined.
+Implementation details.
+The reported results are averaged over 5 runs with different data splits, using test sets comprising 25% of the data and a fixed validation set size of 100, following the methodology of HaloScope Du et al.
+(2024).
+To ensure a fair comparison, we consider two implementations of the latter: a standard setting with $20$ generations (see Tables 5-6), and an efficiency-comparable one with minimal number of generations — $1$ for SelfCheckGPT, $5$ for semantic entropy and EigenScore (Tables 1-2).
+Appendix F provides additional information on implementation details.
+Table 1: ROC AUC $(\uparrow)$ of hallucination detection techniques for three LLMs. The best results for each model are highlighted in bold, and the second best are underlined.
 Method Single MS MARCO CNN/DM + CoQA SQuAD XSum
@@
 Mistral-7B
-SelfCheckGPT [1] 0.62 $\pm$ 0.03 0.63 $\pm$ 0.04 0.93 $\pm$ 0.02 0.82 $\pm$ 0.03 0.71 $\pm$ 0.03
-Semantic entropy [2] 0.54 $\pm$ 0.06 0.57 $\pm$ 0.04 0.84 $\pm$ 0.02 0.71 $\pm$ 0.05 0.65 $\pm$ 0.05
-EigenScore [3] 0.58 $\pm$ 0.02 0.55 $\pm$ 0.04 0.73 $\pm$ 0.02 0.54 $\pm$ 0.04 0.57 $\pm$ 0.03
-Haloscope [4] 0.56 $\pm$ 0.05 0.54 $\pm$ 0.05 0.84 $\pm$ 0.02 0.97 $\pm$ 0.03 0.59 $\pm$ 0.05
-LLM-Check [5] 0.5 $\pm$ 0.06 0.57 $\pm$ 0.03 0.62 $\pm$ 0.04 0.56 $\pm$ 0.06 0.57 $\pm$ 0.03
-Perplexity [6] 0.67 $\pm$ 0.03 0.62 $\pm$ 0.04 0.77 $\pm$ 0.01 0.45 $\pm$ 0.06 0.66 $\pm$ 0.03
-Max entropy [7] 0.66 $\pm$ 0.05 0.6 $\pm$ 0.05 0.73 $\pm$ 0.02 0.73 $\pm$ 0.04 0.71 $\pm$ 0.03
-TOHA (ours) 0.76 $\pm$ 0.04 0.63 $\pm$ 0.04 0.89 $\pm$ 0.02 0.84 $\pm$ 0.01 0.61 $\pm$ 0.03
+SelfCheckGPT [1] 0.63 $\pm$ 0.04 0.51 $\pm$ 0.04 0.86 $\pm$ 0.02 0.71 $\pm$ 0.04 0.66 $\pm$ 0.04
+Semantic entropy [2] 0.54 $\pm$ 0.03 0.51 $\pm$ 0.04 0.83 $\pm$ 0.02 0.70 $\pm$ 0.03 0.56 $\pm$ 0.03
+EigenScore [3] 0.54 $\pm$ 0.04 0.50 $\pm$ 0.06 0.74 $\pm$ 0.02 0.71 $\pm$ 0.04 0.58 $\pm$ 0.04
+HaloScope [4] 0.57 $\pm$ 0.08 0.51 $\pm$ 0.10 0.62 $\pm$ 0.08 0.92 $\pm$ 0.07 0.62 $\pm$ 0.02
+LLM-Check [5] 0.49 $\pm$ 0.03 0.49 $\pm$ 0.03 0.60 $\pm$ 0.01 0.50 $\pm$ 0.03 0.58 $\pm$ 0.04
+Perplexity [6] 0.45 $\pm$ 0.01 0.54 $\pm$ 0.02 0.54 $\pm$ 0.03 0.81 $\pm$ 0.05 0.54 $\pm$ 0.06
+Max entropy [7] 0.68 $\pm$ 0.04 0.60 $\pm$ 0.07 0.73 $\pm$ 0.00 0.75 $\pm$ 0.05 0.71 $\pm$ 0.02
+ReDEEP [8] 0.54 $\pm$ 0.02 0.47 $\pm$ 0.06 0.59 $\pm$ 0.03 0.45 $\pm$ 0.05 0.63 $\pm$ 0.04
+TOHA (ours) 0.76 $\pm$ 0.04 0.60 $\pm$ 0.09 0.89 $\pm$ 0.01 0.96 $\pm$ 0.01 0.66 $\pm$ 0.05
 LLama-2-7B
-SelfCheckGPT [1] 0.6 $\pm$ 0.02 0.62 $\pm$ 0.05 0.78 $\pm$ 0.04 0.59 $\pm$ 0.05 0.70 $\pm$ 0.06
-Semantic entropy [2] 0.54 $\pm$ 0.05 0.54 $\pm$ 0.04 0.76 $\pm$ 0.03 0.63 $\pm$ 0.06 0.63 $\pm$ 0.05
-EigenScore [3] 0.54 $\pm$ 0.02 0.52 $\pm$ 0.04 0.68 $\pm$ 0.03 0.53 $\pm$ 0.05 0.63 $\pm$ 0.06
-Haloscope [4] 0.53 $\pm$ 0.06 0.48 $\pm$ 0.02 0.74 $\pm$ 0.04 0.58 $\pm$ 0.04 0.58 $\pm$ 0.08
-LLM-Check [5] 0.46 $\pm$ 0.02 0.49 $\pm$ 0.03 0.6 $\pm$ 0.03 0.58 $\pm$ 0.05 0.58 $\pm$ 0.08
-Perplexity [6] 0.66 $\pm$ 0.01 0.57 $\pm$ 0.04 0.73 $\pm$ 0.03 0.57 $\pm$ 0.09 0.58 $\pm$ 0.05
-Max entropy [7] 0.66 $\pm$ 0.03 0.57 $\pm$ 0.05 0.71 $\pm$ 0.04 0.64 $\pm$ 0.02 0.56 $\pm$ 0.05
-TOHA (ours) 0.67 $\pm$ 0.02 0.57 $\pm$ 0.05 0.88 $\pm$ 0.04 0.9 $\pm$ 0.04 0.63 $\pm$ 0.02
+SelfCheckGPT [1] 0.59 $\pm$ 0.03 0.60 $\pm$ 0.03 0.66 $\pm$ 0.03 0.57 $\pm$ 0.03 0.64 $\pm$ 0.05
+Semantic entropy [2] 0.53 $\pm$ 0.03 0.51 $\pm$ 0.03 0.76 $\pm$ 0.01 0.73 $\pm$ 0.03 0.61 $\pm$ 0.04
+EigenScore [3] 0.55 $\pm$ 0.03 0.53 $\pm$ 0.04 0.61 $\pm$ 0.03 0.75 $\pm$ 0.02 0.63 $\pm$ 0.02
+HaloScope [4] 0.51 $\pm$ 0.05 0.48 $\pm$ 0.05 0.61 $\pm$ 0.04 0.67 $\pm$ 0.04 0.57 $\pm$ 0.07
+LLM-Check [5] 0.44 $\pm$ 0.02 0.49 $\pm$ 0.06 0.60 $\pm$ 0.03 0.49 $\pm$ 0.01 0.61 $\pm$ 0.01
+Perplexity [6] 0.54 $\pm$ 0.04 0.44 $\pm$ 0.03 0.74 $\pm$ 0.02 0.46 $\pm$ 0.03 0.56 $\pm$ 0.09
+Max entropy [7] 0.65 $\pm$ 0.04 0.59 $\pm$ 0.06 0.65 $\pm$ 0.03 0.73 $\pm$ 0.04 0.56 $\pm$ 0.03
+ReDEEP [8] 0.54 $\pm$ 0.04 0.52 $\pm$ 0.04 0.72 $\pm$ 0.04 0.42 $\pm$ 0.08 0.54 $\pm$ 0.06
+TOHA (ours) 0.65 $\pm$ 0.02 0.56 $\pm$ 0.02 0.90 $\pm$ 0.01 0.87 $\pm$ 0.04 0.68 $\pm$ 0.05
 LLaMA-2-13B
-SelfCheckGPT [1] 0.57 $\pm$ 0.04 0.60 $\pm$ 0.04 0.86 $\pm$ 0.04 0.78 $\pm$ 0.03 0.61 $\pm$ 0.05
-Semantic entropy [2] 0.61 $\pm$ 0.03 0.51 $\pm$ 0.05 0.75 $\pm$ 0.05 0.72 $\pm$ 0.02 0.64 $\pm$ 0.04
-EigenScore [3] 0.56 $\pm$ 0.04 0.49 $\pm$ 0.04 0.45 $\pm$ 0.04 0.51 $\pm$ 0.04 0.55 $\pm$ 0.02
-Haloscope [4] 0.55 $\pm$ 0.05 0.50 $\pm$ 0.01 0.65 $\pm$ 0.04 0.54 $\pm$ 0.03 0.59 $\pm$ 0.01
-LLM-Check [5] 0.44 $\pm$ 0.04 0.57 $\pm$ 0.04 0.54 $\pm$ 0.03 0.53 $\pm$ 0.05 0.63 $\pm$ 0.04
-Perplexity [6] 0.65 $\pm$ 0.02 0.59 $\pm$ 0.02 0.58 $\pm$ 0.04 0.5 $\pm$ 0.05 0.58 $\pm$ 0.05
-Max entropy [7] 0.58 $\pm$ 0.03 0.55 $\pm$ 0.06 0.70 $\pm$ 0.06 0.78 $\pm$ 0.04 0.56 $\pm$ 0.04
-TOHA (ours) 0.69 $\pm$ 0.01 0.54 $\pm$ 0.04 0.93 $\pm$ 0.03 0.92 $\pm$ 0.02 0.66 $\pm$ 0.05
-Table 2: ROC AUC ()($\uparrow$)( ↑ ) of hallucination detection techniques. The best results for each model are highlighted in bold, and the second best are underlined.
-Method Single gen. SQuAD CoQA XSum
-LLaMA-3.1-8B
-SelfCheckGPT [1] 0.79 $\pm$ 0.05 0.76 $\pm$ 0.07 0.81 $\pm$ 0.02
-Semantic entropy [2] 0.58 $\pm$ 0.04 0.83 $\pm$ 0.05 0.49 $\pm$ 0.05
-EigenScore [3] 0.52 $\pm$ 0.07 0.82 $\pm$ 0.06 0.49 $\pm$ 0.05
-Haloscope [4] 0.85 $\pm$ 0.02 0.56 $\pm$ 0.07 0.55 $\pm$ 0.04
-LLM-Check [5] 0.48 $\pm$ 0.04 0.5 $\pm$ 0.08 0.56 $\pm$ 0.03
-Perplexity [6] 0.82 $\pm$ 0.02 0.69 $\pm$ 0.04 0.61 $\pm$ 0.04
-Max entropy [7] 0.5$\pm$ 0.04 0.53 $\pm$ 0.05 0.47 $\pm$ 0.01
-TOHA (ours) 0.85 $\pm$ 0.02 0.73 $\pm$ 0.01 0.63 $\pm$ 0.04
-Qwen2.5-7B
-SelfCheckGPT [1] 0.62 $\pm$ 0.06 0.84 $\pm$ 0.04 0.74 $\pm$ 0.04
-Semantic entropy [2] 0.68 $\pm$ 0.03 0.77 $\pm$ 0.06 0.65 $\pm$ 0.04
-EigenScore [3] 0.67 $\pm$ 0.04 0.66 $\pm$ 0.09 0.51 $\pm$ 0.04
-Haloscope [4] 0.59 $\pm$ 0.05 0.71 $\pm$ 0.05 0.58 $\pm$ 0.05
-LLM-Check [5] 0.5 $\pm$ 0.05 0.54 $\pm$ 0.1 0.55 $\pm$ 0.05
-Perplexity [6] 0.63 $\pm$ 0.06 0.67 $\pm$ 0.08 0.65 $\pm$ 0.05
-Max entropy [7] 0.73 $\pm$ 0.05 0.74 $\pm$ 0.05 0.52 $\pm$ 0.08
-TOHA (ours) 0.8 $\pm$ 0.02 0.69 $\pm$ 0.05 0.69 $\pm$ 0.02
+SelfCheckGPT [1] 0.58 $\pm$ 0.04 0.58 $\pm$ 0.05 0.77 $\pm$ 0.02 0.64 $\pm$ 0.03 0.60 $\pm$ 0.04
+Semantic entropy [2] 0.57 $\pm$ 0.04 0.54 $\pm$ 0.03 0.76 $\pm$ 0.04 0.65 $\pm$ 0.03 0.60 $\pm$ 0.03
+EigenScore [3] 0.56 $\pm$ 0.04 0.47 $\pm$ 0.04 0.57 $\pm$ 0.03 0.57 $\pm$ 0.02 0.52 $\pm$ 0.06
+HaloScope [4] 0.54 $\pm$ 0.09 0.51 $\pm$ 0.04 0.57 $\pm$ 0.03 0.55 $\pm$ 0.02 0.55 $\pm$ 0.07
+LLM-Check [5] 0.49 $\pm$ 0.06 0.56 $\pm$ 0.05 0.57 $\pm$ 0.02 0.57 $\pm$ 0.07 0.57 $\pm$ 0.07
+Perplexity [6] 0.54 $\pm$ 0.04 0.46 $\pm$ 0.07 0.62 $\pm$ 0.03 0.45 $\pm$ 0.02 0.49 $\pm$ 0.05
+Max entropy [7] 0.62 $\pm$ 0.03 0.53 $\pm$ 0.06 0.66 $\pm$ 0.03 0.78 $\pm$ 0.02 0.59 $\pm$ 0.04
+ReDEEP [8] 0.62 $\pm$ 0.06 0.48 $\pm$ 0.05 0.73 $\pm$ 0.02 0.48 $\pm$ 0.07 0.58 $\pm$ 0.08
+TOHA (ours) 0.67 $\pm$ 0.04 0.56 $\pm$ 0.05 0.92 $\pm$ 0.02 0.88 $\pm$ 0.05 0.66 $\pm$ 0.03
+4.2 Results
 Main results.
 The results of our experiments are provided in Tables 12.
-We evaluate TOHA against state-of-the-art hallucination detection methods and demonstrate its competitive performance, consistently securing first or second place across most benchmark datasets.
-TOHA significantly outperforms uncertainty-based baselines and matches the accuracy of consistency-based approaches, achieving a notable $13.4\%$ improvement on the challenging MS MARCO dataset that includes long and detailed model responses.
-While SelfCheckGPT emerges as TOHA’s closest competitor, it relies on additional generations, incurring substantial computational overhead.
-Comparison of the best-performing baselines (Figure 6) confirms that TOHA reduces inference time by an order of magnitude compared to methods of similar quality.
-Compared to Haloscope Du et al.
-(2024), which operates with limited annotated data and large-scale unlabeled data, TOHA not only delivers superior performance but also eliminates the need for unannotated generations for hyperparameter tuning and a separate classifier training.
-This makes TOHA less data-dependent and more practical for real-world applications.
-Another interesting comparison is with LLM-Check Sriramanan et al.
-(2024), which also uses attention maps to compute hallucination scores—specifically, by averaging the log determinant of attention maps from a single pre-selected layer.
-However, our TOHA achieves superior performance, demonstrating that not all attention heads contribute equally to hallucination detection.
-By employing a topology-based head selection strategy, we significantly enhance detection quality.
-Generalizability to different data distributions.
-From a deployment perspective, hallucination detection methods must remain robust to shifts in input data distribution, given the inherent diversity of real-world user queries.
-To evaluate TOHA’s robustness in this regard, we conducted transfer experiments on Mistral-7B (see Figure 4a).
-The results highlight TOHA’s strong transferability: for the XSum and CNN/DM datasets, performance changes in transfer settings fall within the standard deviation.
-For the remaining datasets, TOHA maintains competitive performance compared to baseline methods (Table 1), demonstrating its adaptability to diverse data distributions.
+We evaluate TOHA against state-of-the-art hallucination detection methods and demonstrate its competitive performance TOHA significantly outperforms uncertainty-based baselines and matches the quality of consistency-based approaches, achieving notable improvements of $11.7\%$ on the challenging long-form QA MS MARCO dataset for Mistral-7B and $21.6\%$ on the conversational QA dataset CoQA for LLama-2-7B.
+To evaluate TOHA’s robustness to data distribution change, we conducted transfer experiments on Mistral-7B, see Figure 3(c) for the results.
+TOHA possesses strong transferability: for the XSum and CNN/DM datasets, performance in transfer settings fall within the standard deviation of the method (see Table 9), while remaining competitive on the other datasets (Table 1).
+Evaluation on a multi-hop dataset.
+To validate TOHA in a more realistic setting, we consider an additional experiment using the HotpotQA Yang et al.
+(2018) dataset.
+It consists of questions that require knowledge from multiple supporting documents — much like real-world queries, which also rarely rely on a single source of information.
+The results are provided in Table 3.
+TOHA demonstrates superior performance to baselines, confirming its effectiveness “in the wild”.
+Efficiency.
+Figure 4(a) shows that TOHA is approximately seven times faster than SelfCheckGPT with a single additional generation.
+Given that SelfCheckGPT typically requires 10-20 generations, this makes TOHA over 70 times faster in practice.
+Furthermore, TOHA’s runtime is close to the lightweight entropy baseline; with a more efficient low-level implementation, it has the potential to approach the cost of a single forward pass while achieving higher accuracy than other inexpensive alternatives.
+Table 2: ROC AUC $(\uparrow)$ of hallucination detection techniques. The best results for each model are highlighted in bold, and the second best are underlined.
+Method Single gen. CoQA SQuAD XSum
+LLaMA-3.1-8B
+SelfCheckGPT [1] 0.91 $\pm$ 0.01 0.65 $\pm$ 0.05 0.68 $\pm$ 0.05
+Semantic entropy [2] 0.78 $\pm$ 0.02 0.55 $\pm$ 0.07 0.47 $\pm$ 0.04
+EigenScore [3] 0.80 $\pm$ 0.02 0.45 $\pm$ 0.05 0.50 $\pm$ 0.03
+HaloScope [4] 0.67 $\pm$ 0.08 0.84 $\pm$ 0.07 0.55 $\pm$ 0.06
+LLM-Check [5] 0.54 $\pm$ 0.06 0.49 $\pm$ 0.05 0.52 $\pm$ 0.04
+Perplexity [6] 0.51 $\pm$ 0.05 0.68 $\pm$ 0.02 0.65 $\pm$ 0.03
+Max entropy [7] 0.82 $\pm$ 0.03 0.60 $\pm$ 0.02 0.65 $\pm$ 0.03
+ReDEEP [8] 0.58 $\pm$ 0.08 0.39 $\pm$ 0.04 0.62 $\pm$ 0.06
+TOHA (ours) 0.84 $\pm$ 0.01 0.87 $\pm$ 0.03 0.65 $\pm$ 0.05
+Qwen2.5-7B
+SelfCheckGPT [1] 0.69 $\pm$ 0.01 0.74 $\pm$ 0.02 0.69 $\pm$ 0.02
+Semantic entropy [2] 0.76 $\pm$ 0.04 0.57 $\pm$ 0.04 0.64 $\pm$ 0.03
+EigenScore [3] 0.78 $\pm$ 0.03 0.55 $\pm$ 0.04 0.55 $\pm$ 0.04
+HaloScope [4] 0.66 $\pm$ 0.13 0.75 $\pm$ 0.04 0.57 $\pm$ 0.07
+LLM-Check [5] 0.53 $\pm$ 0.07 0.54 $\pm$ 0.06 0.54 $\pm$ 0.02
+Perplexity [6] 0.39 $\pm$ 0.03 0.65 $\pm$ 0.03 0.53 $\pm$ 0.05
+Max entropy [7] 0.85 $\pm$ 0.02 0.47 $\pm$ 0.05 0.60 $\pm$ 0.06
+ReDEEP [8] 0.37 $\pm$ 0.10 0.56 $\pm$ 0.03 0.67 $\pm$ 0.03
+TOHA (ours) 0.79 $\pm$ 0.05 0.77 $\pm$ 0.02 0.69 $\pm$ 0.03
+Figure 3: (a)-(b): Detection quality dependence on the size of a probe set, models: Mistral-7B (left), LLama-2-7B (right).
+(c) Generalizability between the datasets, model: Mistral-7B.
+The vertical axis corresponds to the origin of the probe set, the horizontal axis to the test dataset.
+Method Single gen. Mistral-7B Llama-2-13B
+SelfCheckGPT 0.70 $\pm$ 0.06 0.63 $\pm$ 0.04
+Semantic entropy 0.70 $\pm$ 0.05 0.70 $\pm$ 0.04
+EigenScore 0.68 $\pm$ 0.04 0.67 $\pm$ 0.04
+HaloScope 0.60 $\pm$ 0.06 0.50 $\pm$ 0.01
+LLM-Check 0.48 $\pm$ 0.03 0.56 $\pm$ 0.04
+Perplexity 0.55 $\pm$ 0.06 0.49 $\pm$ 0.04
+Max entropy 0.62 $\pm$ 0.04 0.69 $\pm$ 0.05
+ReDEEP 0.49 $\pm$ 0.04 0.62 $\pm$ 0.05
+TOHA (ours) 0.71 $\pm$ 0.08 0.80 $\pm$ 0.03
+Table 3: ROC-AUC $(\uparrow)$ values on the HotpotQA dataset. Best results are highlighted in bold, and the second best are underlined.
+4.3 Analysis of Hallucination-Aware Attention Heads
+Attention patterns and copying behaviour.
+For hallucination-aware heads, we analyzed minimal spanning forest (MSF) patterns distinguishing hallucinated and grounded samples.
+A key finding is that for these heads, hallucinated samples frequently exhibit strong attention to the first token, whereas grounded samples tend to attend to the first token less.
+An example is provided in Figure 5.
+Since this pattern is a known behavior of token induction, or copying heads — which default to the first token when unable to find previous occurrences of the current token in the context Elhage et al.
+(2021) — we decided to explore the relationship between these special heads and hallucination-aware ones.
+Figure 4: a): Inference time comparison (seconds) for various methods evaluated on 16 MS MARCO samples using Mistral-7B.
+SelfCheckGPT measurement includes one additional generated answer per sample.
+b)-c): ROC-AUC performance of TOHA across different numbers of selected attention heads ($N_{max}$) on Mistral-7B and Llama-2-7B.
+Figure 5: Attention to the first token (<s> in this example) for (a) a hallucinated generation and (b) a grounded one.
+Green highlights edges and nodes corresponding to grounded tokens, while yellow indicates hallucinated tokens.
+Model: Mistral-7B.
+Figure 6: Copying ranks of hallucination-aware attention heads (lower ranks indicate stronger copying behaviour). The row labels $(X,Y)$ correspond to $X$-th layer and $Y$-th head.
+To investigate this link, we ranked all attention heads in Llama-2-7B and Mistral-7B based on their copying scores, following the method of Feucht et al.
+(2025).
+For the subset of heads frequently identified by TOHA (appearing in $\geq 20\%$ of runs across all datasets), their copying ranks were recorded.
+The results, shown in Figure 6, reveal that hallucination-aware heads are often also among the model’s top-25 copiers, a finding that aligns with prior work Sun et al.
+(2025).
+4.4 Ablation Studies
+Why zero out the distances between prompt tokens?
+We hypothesize that the connections within the prompt contribute little to hallucination detection; therefore, our method is designed to disregard them.
+To validate this architectural choice, we considered an ablation study, where $\operatorname{MTop-Div}_{G}(R,P)$ was replaced by the MST length of the complete graph in Algorithm 1.
+The results showing the effectiveness of the proposed approach are in Table 4.
+Table 4: ROC-AUC $(\uparrow)$ of Algorithm 1 with MST length of the complete graph vs $\operatorname{MTop-Div}_{G}(R,P)$. Best results are highlighted in bold.
+Dataset MST length $\operatorname{MTop-Div}_{G}(R,P)$
+Mistral-7B
+CoQA $0.75\pm 0.03$ $\mathbf{0.90\pm 0.01}$
+MS MARCO $0.38\pm 0.03$ $\mathbf{0.65\pm 0.02}$
+LLama-2-7B
+CoQA $0.60\pm 0.03$ $\mathbf{0.89\pm 0.01}$
+MS MARCO $0.37\pm 0.02$ $\mathbf{0.76\pm 0.04}$
+What about other attention map-based features?
+To demonstrate that our topology-based approach offers an advantage over conventional attention map features, we compared them to $\operatorname{MTop-Div}$ in the supervised setting.
+Table 8 from Appendix B shows that a classifier trained on $\operatorname{MTop-Div}_{G}(R,P)$ achieves the best performance, confirming that our score captures unique patterns in an attention map that standard approaches miss.
 How large should the probe sets be?
-As previously mentioned, TOHA requires only a small set of samples to identify ”hallucination-aware" attention heads.
-To assess its sensitivity to probe set size, we conducted an ablation study (Figure 4).
-The results demonstrate TOHA’s robustness to limited annotated data: even with just $50$ samples, performance does not drop significantly and mostly remains stable as the probe set size increases.
-Figure 4: (a) Generalizability between the datasets, model: Mistral-7B-Instruct.
-The vertical axis corresponds to the origin of the probe set, the horizontal axis  to the test dataset.
-(b)-(c): Detection quality dependence on the size of a probe set, models: Mistral-7B (left), LLama-2-7B (right).
-What do hallucination patterns look like?
-As detailed in Section 3, the topological divergences we employ characterize the MSF connecting the vertices $R$ of the response to the vertices $P$ of the prompt.
-For hallucination-aware heads, we analyzed MSF patterns distinguishing hallucinated and grounded samples.
-A key finding is that hallucinated samples frequently exhibit strong attention to the <s> token, whereas grounded samples tend to attend to <s> less (Figure 5).
-To verify the significance of <s> in hallucination detection, we conducted an ablation experiment: after removing <s> from the texts, we recomputed the TOHA hallucination scores for the selected hallucination-aware heads.
-The results (Table 3) show a significant performance drop, confirming that attention to <s> is a critical indicator of hallucination.
-This finding aligns with prior work demonstrating the influential role of <s> in LLM mechanisms Barbero et al.
+To evaluate TOHA’s sensitivity to probe set size, we conducted a sensitivity study.
+The results from Figure 3 confirm TOHA’s robustness: even with just $50$ samples, the performance drop remains minor with stable results as the probe set size increases.
+How many attention heads do we need?
+To evaluate the sensitivity of our method to the hyperparameter $N_{max}$, we performed an ablation study for values from $1$ to $10$.
+As shown in Figure 4, TOHA achieves strong detection performance even when $N_{\max}=1$, which underscores the effectiveness of our topological approach.
+5 Conclusion
+This paper introduces TOHA, a novel hallucination detection method based on the topological structure of attention maps.
+Central to TOHA is the $\operatorname{MTop-Div}_{G}(R,P)$, our adaptation of Manifold Topology Divergence Barannikov et al.
+(2021) for the graph setting, which, as we demonstrate, serves as a measure of the response novelty in relation to the prompt.
+This property makes the proposed divergence well-suited for hallucination detection in RAG scenarios.
+By analyzing the divergence values, we identified a subset of “hallucination-aware” attention heads that reliably distinguish hallucinated and grounded samples, irrespective of the dataset.
+TOHA computes the final hallucination scores by averaging the topological divergences from these heads.
+Further investigation reveals that some of these heads are associated with copying behavior, which aligns with prior work Sun et al.
 (2025).
-However, using the average attention to <s> alone as a hallucination score proves insufficient (see Table 9 in Appendix D).
-In contrast, our proposed score, which incorporates the intricate structure of attention maps, demonstrates far greater discriminative power for this task.
-Table 3: TOHA performance with and without <s> token. TOP-1 results are highlighted with bold font.
-Dataset with <s> w/o <s>
-Mistral-7B
-MS MARCO 0.67 $\pm$ 0.02 0.56 $\pm$ 0.02
-CoQA 0.88 $\pm$ 0.02 0.32 $\pm$ 0.03
-LLaMA-2-7B
-MS MARCO 0.76 $\pm$ 0.04 0.66 $\pm$ 0.03
-CoQA 0.89 $\pm$ 0.02 0.56 $\pm$ 0.04
-Figure 5: Attention to <s>: a) hallucinated sample and b) grounded one.
-Green color denotes edges and nodes that correspond to grounded tokens of a generation, yellow color — hallucinated ones.
-Model: Mistral-7B.
-Figure 6: Comparison of methods’ inference time in seconds.
-The measurements were obtained for $16$ random samples from MS MARCO, model: Mistral-7B.
-For the SelfCheckGPT, $20$ additional answers were generated for each sample.
-5 Related works
+Extensive experiments show that TOHA is a robust alternative to existing approaches, matching or surpassing state-of-the-art baselines.
+Moreover, our method is both data- and compute-efficient: just $50$ annotated samples suffice for reliable detection, and inference runs several times faster than comparable methods of similar quality.
+Crucially, we validate TOHA’s transferability, demonstrating its robustness to shifts in data distribution — a key advantage for real-world deployment, where LLM inputs are far more diverse and complex than benchmark examples.
+In summary, TOHA delivers state-of-the-art detection performance while combining efficiency and solid generalizability, making it particularly well-suited for practical applications.
+Limitations
+While TOHA demonstrates strong performance and efficiency, several limitations warrant discussion.
+Model-specific dependencies.
+TOHA’s effectiveness relies on identifying “hallucination-aware” attention heads, which may vary across LLM architectures.
+While our experiments cover popular open-source models (e.g., LLaMA, Mistral), further validation is needed for proprietary or larger models (e.g., GPT-4, Claude).
+Multimodal extensions.
+The current framework operates solely on text.
+Adapting TOHA to multimodal settings (e.g., vision-language models) would require redefining attention graphs across heterogeneous data modalities.
+Appendix A Related Works
 Hallucination detection methods.
@@
 (2024); Chen et al.
-(2024); Kuhn et al.
 (2024); Qiu and Miikkulainen (2024); Nikitin et al.
+(2024); Hou et al.
 (2025) that use the diversity of multiple LLM responses as a hallucination score offer robust detection but impose significant computational overhead.
 Surface-level techniques like perplexity and logit entropy Fadeeva et al.
-(2024); Malinin and Gales (2021) analyze model confidence directly from output distributions — efficient but limited in detection capability as they neglect the model’s rich internal representations.
+(2024); Malinin and Gales (2021) analyze model confidence directly from output distributions — efficient but limited in detection capability as they neglect the model’s rich internal representations Gekhman et al.
+(2025).
 Hidden states-based classifiers Azaria and Mitchell (2023); Sky et al.
@@
 (2025) require extensive annotated datasets, which are scarce in the public domain Zhang et al.
+(2023), and suffer from poor task transferability Sky et al.
 (2024).
-This issue was partially addressed by the Haloscope Du et al.
-(2024), which leverages unlabeled data “in the wild” with minimum annotated data needed for hyperparameter selection.
+While HaloScope Du et al.
+(2024) partially addresses the data scarcity issue by leveraging unlabeled “in-the-wild” generations, it still demands a large volume of unannotated model outputs to achieve strong performance.
 Attention map-based methods represent a promising yet underdeveloped direction.
-Current techniques either rely on large labeled data, e.g., Lookback Lens Chuang et al.
-(2024), or exploit only simple attention graph properties, such as self-loop weights in LLM-Check Sriramanan et al.
-(2024).
-This leaves a critical research gap: training-free methods that fully leverage the rich structural information encoded in attention relationships remain underexplored.
+Current techniques either (i) rely on large annotated datasets (e.g., Lookback Lens Chuang et al.
+(2024), LapEigVals Binkowski et al.
+(2025)), (ii) exploit only simplistic attention graph properties like self-loop weights (LLM-Check Sriramanan et al.
+(2024)), or (iii) disregard attention map geometry entirely, using mechanistic scores instead (ReDEEP Sun et al.
+(2025)).
+This leaves a critical research gap: training-free methods that fully leverage the rich structural information encoded in attention graphs remain underexplored.
 Topological Data Analysis (TDA) in NLP.
@@
 (2023) and grammatical acceptability classification Proskurina et al.
-(2023b), where topological features extracted from the attention graphs were used as input to train auxiliary classifiers.
-6 Conclusion
-This paper introduces TOHA, a novel hallucination detection method based on the topological divergence of attention maps.
-At its core, TOHA leverages our key observation that specific attention heads exhibit consistent patterns during hallucinations — regardless of the dataset.
-TOHA computes hallucination scores by averaging the topological divergences from these heads, and we formally prove several stability properties to ensure these scores are reliable.
-Additionally, we explored the behaviour of “hallucination-aware” heads, discovering that the attention to <s> token plays an important role in their discriminative ability.
-This importance of <s> aligns well with prior work Barbero et al.
+(2023), where topological features extracted from the attention graphs were used as input to train auxiliary classifiers.
+Appendix B Additional Experiment Results
+B.1 Multiple generations-based methods
+To provide a complete comparison, we considered an implementation of consistency-based methods with $20$ additional generations.
+The results are demonstrated in Tables 5-6.
+We can see that TOHA still remains superior to the baselines, achieving top performance in most experiments.
+Table 5: ROC AUC $(\uparrow)$ of multiple generations-based methods (with 20 additional samples) and TOHA. The best results for each model are highlighted in bold, and the second best are underlined.
+Method Single MS MARCO CNN/DM + CoQA SQuAD XSum
+generation Recent News
+Mistral-7B
+SelfCheckGPT [1] 0.67 $\pm$ 0.03 0.59 $\pm$ 0.04 0.93 $\pm$ 0.01 0.83 $\pm$ 0.02 0.71 $\pm$ 0.04
+Semantic entropy [2] 0.53 $\pm$ 0.03 0.52 $\pm$ 0.03 0.86 $\pm$ 0.01 0.74 $\pm$ 0.01 0.63 $\pm$ 0.02
+EigenScore [3] 0.49 $\pm$ 0.03 0.53 $\pm$ 0.05 0.78 $\pm$ 0.02 0.77 $\pm$ 0.04 0.59 $\pm$ 0.05
+TOHA (ours) 0.76 $\pm$ 0.04 0.60 $\pm$ 0.09 0.89 $\pm$ 0.01 0.96 $\pm$ 0.01 0.66 $\pm$ 0.05
+LLama-2-7B
+SelfCheckGPT [1] 0.60 $\pm$ 0.04 0.60 $\pm$ 0.03 0.77 $\pm$ 0.02 0.78 $\pm$ 0.02 0.67 $\pm$ 0.04
+Semantic entropy [2] 0.56 $\pm$ 0.03 0.49 $\pm$ 0.03 0.79 $\pm$ 0.01 0.77 $\pm$ 0.02 0.63 $\pm$ 0.04
+EigenScore [3] 0.57 $\pm$ 0.04 0.52 $\pm$ 0.06 0.61 $\pm$ 0.03 0.78 $\pm$ 0.02 0.65 $\pm$ 0.03
+TOHA (ours) 0.65 $\pm$ 0.02 0.56 $\pm$ 0.02 0.90 $\pm$ 0.01 0.87 $\pm$ 0.04 0.68 $\pm$ 0.05
+LLaMA-2-13B
+SelfCheckGPT [1] 0.61 $\pm$ 0.05 0.60 $\pm$ 0.06 0.88 $\pm$ 0.02 0.75 $\pm$ 0.04 0.61 $\pm$ 0.04
+Semantic entropy [2] 0.60 $\pm$ 0.03 0.52 $\pm$ 0.03 0.77 $\pm$ 0.04 0.70 $\pm$ 0.02 0.62 $\pm$ 0.03
+EigenScore [3] 0.58 $\pm$ 0.04 0.48 $\pm$ 0.05 0.59 $\pm$ 0.03 0.60 $\pm$ 0.02 0.54 $\pm$ 0.05
+TOHA (ours) 0.67 $\pm$ 0.04 0.56 $\pm$ 0.05 0.92 $\pm$ 0.02 0.88 $\pm$ 0.05 0.66 $\pm$ 0.03
+Table 6: ROC AUC $(\uparrow)$ of multiple generations-based methods (with 20 additional samples) and TOHA. The best results for each model are highlighted in bold, and the second best are underlined.
+Method Single gen. CoQA SQuAD XSum
+LLaMA-3.1-8B
+SelfCheckGPT [1] 0.95 $\pm$ 0.01 0.78 $\pm$ 0.03 0.75 $\pm$ 0.03
+Semantic entropy [2] 0.82 $\pm$ 0.03 0.54 $\pm$ 0.06 0.46 $\pm$ 0.05
+EigenScore [3] 0.84 $\pm$ 0.01 0.56 $\pm$ 0.03 0.48 $\pm$ 0.03
+TOHA (ours) 0.84 $\pm$ 0.01 0.87 $\pm$ 0.03 0.65 $\pm$ 0.05
+Qwen2.5-7B
+SelfCheckGPT [1] 0.75 $\pm$ 0.02 0.77 $\pm$ 0.02 0.72 $\pm$ 0.03
+Semantic entropy [2] 0.83 $\pm$ 0.05 0.58 $\pm$ 0.04 0.69 $\pm$ 0.05
+EigenScore [3] 0.83 $\pm$ 0.05 0.56 $\pm$ 0.03 0.48 $\pm$ 0.03
+TOHA (ours) 0.79 $\pm$ 0.05 0.77 $\pm$ 0.02 0.69 $\pm$ 0.03
+B.2 Comparison to ReDEEP
+For a comprehensive comparison with ReDEEP Sun et al.
+(2025), which was evaluated on the entire RAGTruth dataset in the original paper, we conducted this evaluation of TOHA under the same conditions.
+As shown in Table 7, TOHA not only outperforms ReDEEP but also demonstrates greater robustness to data distribution shifts, exhibiting less significant performance degradation on separate benchmarks of RAGTruth (Tables 1-2).
+Method Llama-2-7B Llama-2-13B
+ReDEEP [8] 0.68 $\pm$ 0.02 0.77 $\pm$ 0.01
+TOHA (ours) 0.70 $\pm$ 0.04 0.80 $\pm$ 0.02
+Table 7: Performance comparison between ReDEEP and TOHA on the entire RAGTruth dataset.
+B.3 Alternative attention map-based features for hallucination detection
+Table 8: ROC-AUC values of supervised classifiers on top of various set of features. TOP-1 results are highlighted with bold font, while TOP-2 are underlined.
+Features MS MARCO CoQA
+Mistral-7B
+Standard topological 0.67 0.69
+Sparsity ratio 0.66 0.7
+Entropy 0.75 0.77
+Wasserstein 0.77 0.73
+Spectral norm 0.73 0.72
+Attention to <s> 0.65 0.61
+MTop-Div 0.86 0.98
+LLaMA-2-7B
+Standard topological 0.69 0.7
+Sparsity ratio 0.49 0.61
+Entropy 0.38 0.68
+Wasserstein 0.73 0.6
+Spectral norm 0.49 0.64
+Attention to <s> 0.62 0.64
+MTop-Div 0.75 0.96
+During preliminary experiments for an attention map-based hallucination detector, we evaluated a range of topological and traditional features.
+As standard topological features, we employed barcode-based features, such as the sum of bar lengths in persistence diagrams, and naive topological features, including the average vertex degree in attention graphs Kushnareva et al.
 (2021).
-Extensive experiments show that TOHA is a robust alternative to existing approaches, matching or surpassing state-of-the-art baselines like SelfCheckGPT Manakul et al.
+As for traditional attention-based features, we used sparsity ratio, attention entropy, and spectral norm Kobayashi et al.
+(2020); Vig and Belinkov (2019); Ji et al.
 (2021).
-Notably, TOHA is significantly more efficient, operating up to an order of magnitude faster than methods of comparable quality.
-We further validate TOHA’s transferability, demonstrating its robustness to shifts in data distribution — a critical advantage for real-world deployment, where LLM inputs are far more diverse and complex than benchmark examples.
-In summary, TOHA delivers state-of-the-art detection performance while combining efficiency and solid generalizability, making it particularly suited for practical applications.
-Limitations
-While TOHA demonstrates strong performance and efficiency, several limitations warrant discussion.
-RAG scenario.
-While TOHA operates effectively in the RAG scenario under the assumption that the provided context contains the correct answer, we recognize that this condition may not always hold in real-world applications.
-This limitation points to an important direction for future research, where the method could be extended to handle cases of incomplete or unreliable context knowledge.
-Model-specific dependencies.
-TOHA’s effectiveness relies on identifying “hallucination-aware” attention heads, which may vary across LLM architectures.
-While our experiments cover popular open-source models (e.g., LLaMA, Mistral), further validation is needed for proprietary or larger models (e.g., GPT-4, Claude).
-Multimodal extensions.
-The current framework operates solely on text.
-Adapting TOHA to multimodal settings (e.g., vision-language models) would require redefining attention graphs across heterogeneous data modalities.
-Appendix A Topological data analysis: background
+We also considered Wasserstein distances between the persistent diagrams Edelsbrunner and Harer (2010) of the prompt and response subgraphs as an alternative way to measure their similarity.
+Finally, we analyzed the average attention to the first token as we discovered that hallucination-aware heads often attend to in presence of hallucinations (see Section 4).
+To identify the most informative features for hallucination detection, we trained $L1$-regularized supervised classifiers on concatenated features from all layers and heads and compared them to classifiers trained only on $\operatorname{MTop-Div}_{G}(R,P)$ values under the same conditions.
+The results are presented in Table 8.
+While the classifier based on $\operatorname{MTop-Div}_{G}(R,P)$ values significantly outperforms alternative approaches, computing these values across all layers and attention heads is highly computationally expensive.
+To address this, we developed TOHA — a more efficient alternative that aggregates divergence values from only a subset of the “hallucination-aware” attention heads.
+B.4 Other metrics for the head selection procedure
+Additionally, we investigated alternative attention-map-based scores — including entropy, spectral norm, and the Wasserstein distance between the persistent diagrams of prompts and responses — for selecting specialized attention heads.
+Following the pipeline of the Algorithm 1, we computed the average distances between hallucinated and grounded samples using alternative scores.
+The results, presented in Figure 10, reveal that the heads that separate samples best for MS MARCO do not generalize to the CoQA dataset.
+This suggests that our proposed $\operatorname{MTop-Div}_{G}(R,P)$ metric is better suited for the task compared to existing solutions.
+B.5 Performance analysis
+Additionally, we analysed TOHA’s performance for extremely short responses consisting of one word.
+The results are presented in Figure 7.
+While its performance drops — as expected when the response graph degenerates to a single vertex — it remains non-random.
+This indicates that TOHA retains some predictive power even for these challenging, out-of-design cases.
+As for the TOHA’s performance on longer responses, the results for the MS MARCO, CNN/DM and XSum datasets demonstrate that as the response length grows, the hallucination signals become less distinguishable for all the considered methods, as hallucinated tokens comprise a very small part of the response.
+However, TOHA remains superior to baselines, confirming that our topology-based approach is more effective than existing methods for complex long-form responses.
+Figure 7: Comparison of ROC-AUC scores ($\uparrow$) for single-word versus multi-word model responses. Dataset: SQuAD.
+B.6 P-values for the transferability experiment
+As we mentioned, for CNN/DM and Xsum dataset, the TOHA’s performance in transfer settings fall within the standard deviation of the method.
+To confirm this, we calculated the p-vales of the corresponding T-test for means of two independent samples.
+The results, provided in Table 9, support our claim: all p-values are significantly greater than $0.05$.
+Probe set CNN/DM XSum
+MS MARCO 0.272 1.0
+CNN/DM 1.0 0.746
+CoQA 0.499 0.134
+SQuAD 0.373 0.346
+XSum 0.515 1.0
+Table 9: P-values of the T-test for the transferability experiment.
+Appendix C Topological Data Analysis: Background
 A simplicial complex $S$ is a collection of simplices such that every face of a simplex $\sigma\in S$ is also in $S$.
@@
 Formally, homology groups are derived from a chain complex $(C_{\bullet},\partial_{\bullet})$, which is a sequence of $C_{k}$ connected by boundary maps $\partial_{k}$:
-$C_{\bullet}:\cdots\rightarrow C_{k+1}\xrightarrow{\partial_{k+1}}C_{k}\xrightarrow{\partial_{k}}\cdots,\,$
+$C_{\bullet}:\cdots\rightarrow C_{k+1}\xrightarrow{\partial_{k+1}}C_{k}\xrightarrow{\partial_{k}}\cdots,\,\$
 $\partial_{k}\circ\partial_{k+1}=0.$
@@
 $\subseteq VR_{\varepsilon_{n}}(X)=VR(X).$
-[…]
+As the threshold $\varepsilon$ increases, new topological features (e.g., connected components, holes) can appear and disappear.
+The persistent homology tool tracks the dynamics of these topological features.
+Formally, the $k$-th persistent homology of $S$ is the pair of sets of vector spaces $\{H_{k}(S_{i})\,|\,0\leq i\leq n\}$ and maps $f_{ij}$, where $f_{ij}:H_{k}(S_{i})\rightarrow H_{k}(S_{j})$ is a map induced by the embedding $S_{i}\subseteq S_{j}$.
+Each persistent homology class in this sequence is “born” at some $S_{i}$ and “dies” at some $S_{j}$ or never dies Barannikov (1994).
+This birth-death process of a basic set of independent topological features can be visualized as the set of intervals $\varepsilon_{\mathrm{birth}},\varepsilon_{\mathrm{death}}]$ called barcode (see Figure 8).
+The features with 0 lifespans are typically excluded.
+The horizontal axis is a sequence of thresholds $\varepsilon$, and each horizontal bar corresponds to a single feature.
+We begin with $|X|=m$ connected components (all of them are “born”), and as $\varepsilon$ increases, their pairs are merged (each merge corresponds to a “death” of a feature).
+The $0$-th barcode construction procedure is equivalent to Kruskal’s algorithm for minimum spanning tree (MST), the bars in the barcode correspond to the edges in the MST of $X$ Tulchinskii et al.
+(2023).
 Figure 8: $H_{0}$ barcode construction.
@@
 The horizontal axis is a sequence of thresholds $\varepsilon$, and each horizontal bar corresponds to a single feature.
-Appendix B MTop-Div on graphs properties
-Basic properties of MTop-Div for attention graphs.
-Now we consider specific properties for our adaptation of $\operatorname{MTop-Div}(R,P)$.
-Proposition B.1.
-The following holds for any attention graph $G$ and its complementary vertex subsets $P,R\subset~{}V_{G}$.
-  • $\operatorname{MTop-Div}(R,P)$ value equals the length of the MSF attaching $R$ to $P$.
-  • Let the natural norm on the cross-barcodes be defined as follows:
-    $\lVert\operatorname{Cross-Barcode}_{0}\rVert_{B}=\\ \max_{[b_{j},d_{j}]\in\operatorname{Cross-Barcode}_{0}}(d_{j}-b_{j}).$ (3)
-    The norm of $\operatorname{Cross-Barcode}_{0}(R,P)$ lays in the interval [0, 1]01[0,\,1][ 0 , 1 ]:
-    $0\leq\lVert\operatorname{Cross-Barcode}_{0}(R,P)\rVert_{B}\leq 1$ (4)
+Appendix D Properties of $\operatorname{MTop-Div}_{G}(R,P)$
+Proof of Proposition 3.1.
+The $0$-th Cross-Barcode coincides with the set of edges in the minimal spanning tree of the weighted graph $G$ with all the weights within $P$ vertex subset equal zero.
+Excluding the zero weight edges, this edge set coincides with the minimal spanning forest attaching the vertex set $R$ to $P$ vertices.
+∎
+Other properties of MTop-Div for attention graphs.
+Here we present some other properties of our proposed $\operatorname{MTop-Div}_{G}(R,P)$.
+Proposition D.1.
+The following holds for any attention graph $G$ with vertex set $V_{G}$ and its complementary vertex subsets $P,R$, where $P\cup R=V_{G}$ and $P\cap R=\varnothing$.
   • The divergence itself is bounded by
-    $0\leq\operatorname{MTop-Div}(R,P)\leq|R|.$ (5)
-The second and third statements are immediately obtained from the properties of an attention matrix: all its weights lie between 00 and $1$.
-The following property formalizes the intuition behind our metric — it measures the strength of the response’s connection to the prompt through multiscale topological features of the attention graph.
-Proposition B.2.
+    $0\leq\operatorname{MTop-Div}_{G}(R,P)\leq|R|.$ (4)
+  • (Stability.) If the weights of $G$ change by no more than $\varepsilon$, then the corresponding $\operatorname{MTop-Div}(R,P)$ changes by no more than $\delta=\varepsilon|R|$.
   • (Exact sequence.) For any $\alpha$, the following sequence of natural maps of homology groups is exact
@@
     $\xrightarrow{r_{1}}H_{0}(VR_{\alpha}(G,w_{(R\cup P)/P}))\xrightarrow{r_{0}}0.$
-Proof of Proposition B.2.
-We have to check the definition of the exact sequence: $\text{Ker}(r_{i})=\text{Im}(r_{i+1})$.
-For a pair $r_{0},\,r_{1}$, it is equivalent to the surjectvity of $r_{1}$.
-The $H_{0}$ homology group of a graph corresponds to the connected components of the graph.
-The set of edges $E_{(G,w)}^{\leq\alpha}=\{e\in E_{G}|w_{e}\leq\alpha\}$ is always a subset in the analogous set of the weighted graph $(G,w_{(R\cup P)/P})$ with all weight edges between $P$ vertices set to zero.
-Therefore, the map $r_{1}$ between their connected components is surjective.
-Similarly, the kernel of the map $r_{1}$ is spanned by the differences of two connected components, which are merged after adding some of the edges between $P$ vertices, and any such difference lies in the image of the map $r_{2}$.
-Also, any two vertices from $P$ belong to the same connected component in the graph $(G,w_{(R\cup P)/P}\leq\alpha)$, hence the image of $r_{2}$ is in the kernel of $r_{1}$.
-Therefore, the considered sequence is exact indeed.
-∎
-Proof of Proposition 3.1.
-1.
-The $0-$th Cross-Barcode coincides with the set of edges in the minimal spanning tree of the weighted graph $G$ with all the weights within $P$ vertex subset equal zero.
-Excluding the zero weight edges, this edge set coincides with the minimal spanning forest attaching the vertex set $R$ to $P$ vertices.
-∎
-2. Denote by $\operatorname{MSF}(R,P)$ the minimum spanning forest attaching $R$ to $P$. Note that we have properties B.1, so
+  • (Connection with hallucinations.) The normalized divergence value $\frac{1}{|R|}\operatorname{MTop-Div}(R,P)=0$ iff the MSF attaches every response token to a prompt token by a subtree with attention weights =1=1.
+Proof of Proposition D.1.
+1. This property is immediately obtained from the properties of an attention map: all its weights lie between 0 and $1$. ∎
+2. Denote by $\operatorname{MSF}(R,P)$ the minimum spanning forest attaching $R$ to $P$. Note that we have properties D.1, so
 $\operatorname{MTop-Div}(R,P)=\sum_{e\in\operatorname{MSF}(R,P)}w(e).$ (5)
@@
 There are two possibilities: 1) after a change, all MSF edges remain the same, or 2) some edges are replaced with other edges.
-In the first case, it is obvious that the total sum of edge weights changes by no more than $\delta=\varepsilon\cdot\#\text{edges}(\operatorname{MSF}(R,P))=\varepsilon\cdot|R|$.
+In the first case, it is obvious that the total sum of edge weights changes by no more than
+$\delta=\varepsilon\cdot\#\text{edges}(\operatorname{MSF}(R,P))=\varepsilon\cdot|R|.$
 Consider the second case.
@@
 ∎
-3. Follows obviously from the MSF formula for $\operatorname{MTop-Div}(R,P)$ and attention map properties.
-Appendix C Datasets
+3.
+We have to check the definition of the exact sequence: $\text{Ker}(r_{i})=\text{Im}(r_{i+1})$.
+For a pair $r_{0},\,r_{1}$, it is equivalent to the surjectvity of $r_{1}$.
+The $H_{0}$ homology group of a graph corresponds to the connected components of the graph.
+The set of edges $E_{(G,w)}^{\leq\alpha}=\{e\in E_{G}|w_{e}\leq\alpha\}$ is always a subset in the analogous set of the weighted graph $(G,w_{(R\cup P)/P})$ with all weight edges between $P$ vertices set to zero.
+Therefore, the map $r_{1}$ between their connected components is surjective.
+Similarly, the kernel of the map $r_{1}$ is spanned by the differences of two connected components, which are merged after adding some of the edges between $P$ vertices, and any such difference lies in the image of the map $r_{2}$.
+Also, any two vertices from $P$ belong to the same connected component in the graph $(G,w_{(R\cup P)/P}\leq\alpha)$, hence the image of $r_{2}$ is in the kernel of $r_{1}$.
+Therefore, the considered sequence is exact indeed.
+∎4.
+Follows obviously from the MSF formula for $\operatorname{MTop-Div}(R,P)$ and attention map properties.
+∎
+Intuition behind MTop-Div properties and hallucination detection.
+The stability property guarantees that similar attention patterns yield similar hallucination scores, making the metric’s behavior consistent and predictable.
+The exact sequence property formalizes geometric intuition behind our metric — it measures the strength of the response’s connection to the prompt through multiscale topological features of the attention graph.
+In the last property, we present an “ideal” case: if a model “knows what to look at” — each token in the response attends to some token in the prompt with an attention weight equal to 1 — $\operatorname{MTop-Div}$ would be equal to 0, indicating zero uncertainty.
+Appendix E Datasets
 SQuAD Rajpurkar et al.
@@
 Table 12: Example of annotation prompt passed to GPT-4o (we add additional delimiter spaces and formatting not present in actual prompts for better readability).
-Prompt number 1 2 3 4 5
-CoQA Accuracy ($\uparrow$) 0.809 ± 0.017 0.861 ± 0.015 0.742 ± 0.003 0.795 ± 0.009 0.831 ± 0.025
-Precision ($\uparrow$) 0.849 ± 0.021 0.911 ± 0.007 0.771 ± 0.003 0.828 ± 0.011 0.860 ± 0.012
-Recall ($\uparrow$) 0.871 ± 0.004 0.877 ± 0.019 0.877 ± 0.013 0.877 ± 0.005 0.893 ± 0.027
-SQuAD Accuracy ($\uparrow$) 0.831 ± 0.003 0.857 ± 0.018 0.857 ± 0.008 0.872 ± 0.003 0.854 ± 0.007
-Precision ($\uparrow$) 0.813 ± 0.002 0.831 ± 0.028 0.845 ± 0.021 0.850 ± 0.011 0.847 ± 0.007
-Recall ($\uparrow$) 0.796 ± 0.008 0.839 ± 0.010 0.823 ± 0.023 0.858 ± 0.018 0.813 ± 0.017
-Average Accuracy ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$)
-CoQA 0.808 0.844 0.879
-SQuAD 0.854 0.837 0.826
-Table 7: Classification metrics of GPT-4o annotation for CoQA and SQuAD with human labels considered actual annotation.
-The top table shows metric scores for different variants of prompts used.
-The bottom table shows the metric scores averaged across all prompt variants.
-C.1 Data Generation & Annotation
+Prompt number 1 2 3 4 5 Average
+CoQA Accuracy ($\uparrow$) 0.809 ± 0.017 0.861 ± 0.015 0.742 ± 0.003 0.795 ± 0.009 0.831 ± 0.025 0.808
+Precision ($\uparrow$) 0.849 ± 0.021 0.911 ± 0.007 0.771 ± 0.003 0.828 ± 0.011 0.860 ± 0.012 0.844
+Recall ($\uparrow$) 0.871 ± 0.004 0.877 ± 0.019 0.877 ± 0.013 0.877 ± 0.005 0.893 ± 0.027 0.879
+SQuAD Accuracy ($\uparrow$) 0.831 ± 0.003 0.857 ± 0.018 0.857 ± 0.008 0.872 ± 0.003 0.854 ± 0.007 0.854
+Precision ($\uparrow$) 0.813 ± 0.002 0.831 ± 0.028 0.845 ± 0.021 0.850 ± 0.011 0.847 ± 0.007 0.837
+Recall ($\uparrow$) 0.796 ± 0.008 0.839 ± 0.010 0.823 ± 0.023 0.858 ± 0.018 0.813 ± 0.017 0.826
+Table 13: Classification metrics of GPT-4o annotation for CoQA and SQuAD with human labels considered as true.
+The table shows metric scores for different variants of prompts used and the score value averaged across all prompt variants.
+E.1 Data Generation & Annotation
 Generation.
@@
 For human annotation, we asked three team members with at least upper-intermediate English proficiency to independently annotate approximately 100 samples from each dataset.
+The human annotation consistency metrics are provided in Table 14.
 We selected samples where all annotators reached a consensus and considered these annotations the ground truth hallucination labels.
+Pair Accuracy Pearson-r
+1 vs 2 0.783 0.6
+1 vs 3 0.812 0.631
+2 vs 3 0.826 0.653
+Mean 0.807 0.628
+Std 0.022 0.027
+Table 14: Cross-annotator consistency metrics.
 To further evaluate GPT-4o, we conducted automatic annotation using several variations of prompts, each reformulating the task for GPT-4o, including zero-shot and few-shot versions.
@@
   1. 1.
-    Rouge-L scoring: we computed Rouge-L scores (using the evaluate library) between the model’s response and the ground-truth answers.
+    Rouge-L scoring: we computed Rouge-L scores (using the evaluate library, v0.4.6) between the model’s response and the ground-truth answers.
   2. 2.
@@
 The procedure outlined above selects a different number of objects in a sample depending on the quality of the model’s responses.
+Additionally, we provide the statistics regarding the response length distribution for each of the datasets we gathered (Figure 9).
 Model CoQA SQuAD XSum
@@
 Mistral-7B 776 776 311 389 301 448
-LLaMA-2-7B 375 375 357 235 239 507
+LLaMA-2-7B 375 375 440 258 239 507
 LLaMA-2-13B 279 384 314 436 208 522
-LLaMA-3.1-8B 189 200 350 400 243 407
-Qwen2.5-7B 124 183 215 249 194 556
+LLaMA-3.1-8B 356 350 350 400 243 407
+Qwen2.5-7B 171 218 423 423 194 556
 Table 15: Datasets statistics. Number of hallucinated and grounded samples of each model.
-Appendix D Other experiment results
-D.1 Alternative attention map-based features for hallucination detection
-Figure 8: $\Delta_{ij}$ values for $ij$-th heads, MS MARCO vs CoQA.
-Vertical axis corresponds to the difference on dataset (B), horizontal — to the one on dataset (A).
-The heads that segregate samples best are highlighted in pink.
-Model names for a row are on the left side.
-Table 9: ROC-AUC values of supervised classifiers on top of various set of features. TOP-1 results are highlighted with bold font, while TOP-2 are underlined.
-Features MS MARCO CoQA
-Mistral-7B
-Standard topological 0.67 0.69
-Sparsity ratio 0.66 0.7
-Entropy 0.75 0.77
-Wasserstein 0.77 0.73
-Spectral norm 0.73 0.72
-Attention to <s> 0.65 0.61
-MTop-Div 0.86 0.98
-LLaMA-2-7B
-Naive topological 0.69 0.7
-Sparsity ratio 0.49 0.61
-Entropy 0.38 0.68
-Wasserstein 0.73 0.6
-Spectral norm 0.49 0.64
-Attention to <s> 0.62 0.64
-MTop-Div 0.75 0.96
-In our preliminary experiments for developing an attention maps-based hallucination detector, we trained classifiers using topological features previously applied to other NLP tasks Kushnareva et al.
-(2021); Cherniavskii et al.
-(2022), as well as traditional attention map characteristics.
-As standard topological features, we considered barcode-based features, such as the sum of bar lengths in persistence diagrams, and naive topological features, including the average vertex degree in attention graphs.
-For traditional attention-based features, we used sparsity ratio, attention entropy, and spectral norm Kobayashi et al.
-(2020); Vig and Belinkov (2019); Ji et al.
-(2021).
-We also considered Wassersein distances between the persistent diagrams Edelsbrunner and Harer (2010) of a context and a response subgraphs as an alternative way to assess their similarity.
-Finally, we analyzed the average attention to <s> token as we discovered that hallucination-aware heads often attend to it when the hallucination is present (see 4).
-To determine the most informative features for hallucination detection, we trained supervised classifiers on concatenated features from all layers and heads and compared them to classifiers trained only on MTop-Div values under the same conditions.
-The results are presented in Table 9.
-While the classifier based on MTop-Div values significantly outperforms alternative approaches, computing these values across all layers and attention heads is highly computationally expensive.
-To address this, we developed TOHA — a more efficient alternative by aggregating MTop-Div values from only a subset of the “hallucination-aware” attention heads in the model.
-D.2 Other metrics for the head selection procedure
-We also investigated alternative attention-map-based scores — including entropy, spectral norm, and the Wasserstein distance between the persistent diagrams of prompts and responses — for selecting specialized attention heads.Following the same pipeline as in TOHA, we computed the average distances between hallucinated and grounded samples using alternative proximity metrics.
-The results, presented in Figure 8, reveal that the most segregating heads for MS MARCO do not generalize to the CoQA dataset.
-This suggests that our proposed MTop-Div metric is better suited for the task compared to existing solutions.
-Appendix E Implementation details
+Figure 9: The distributions of response lengths (in words).
+Figure 10: $\Delta_{ij}$ values for ijij heads, MS MARCO and CoQA.
+Vertical axis corresponds to the difference on the dataset (B), horizontal to the difference on the dataset (A).
+The heads that separate samples best are highlighted in pink.
+Appendix F Implementation details
 In this section, we describe the key implementation choices.
@@
 (2024); Azaria and Mitchell (2023).
-  • For methods that rely on multiple generations, we generated 20 samples per input, following recommendations from Manakul et al. (2024); Chen et al. (2024).
-  • For SelfCheck-GPT, we used its NLI-based variant.
+  • For SelfCheckGPT, we used its NLI-based version.
   • For LLM-Check, we considered its white-box attention score modification, as it works in a setting similar to ours.
-  • For Haloscope, we reserved 150 samples for hyperparameter tuning and treated the rest as unlabeled data.
-For the CoQA dataset, we added several thousand extra generations during training, as in the original paper.
-Other datasets were used as-is, since their sizes are comparable with those in Haloscope’s experiments on the TruthfulQA dataset.
-  • For TOHA’s head selection, we similarly used 150 annotated samples. The topological divergences were calculated using ripser++ library Zhang et al. (2020), MIT license.
-In our experiments, we used 60/15/25 train/val/test split. All the obtained results were averaged over five runs. All experiments were carried out using NVidia L40.
-Appendix F Use of scientific artifacts & AI assistants
+  • The topological divergences were calculated using ripser library Tralie et al. (2018), MIT license.
+  • For the RAGTruth dataset, the model settings were aligned with the original paper.
+All experiments were carried out using NVidia L40.
+Appendix G Use of scientific artifacts & AI assistants
 CoQA contains passages from seven domains under the following licenses: Literature and Wikipedia passages are shared under CC BY-SA 4.0 license; Children’s stories are collected from MCTest which comes with MSR-LA license; Middle/High school exam passages are collected from RACE which comes with its own license; News passages are collected from the DeepMind CNN dataset which comes with Apache license.
@@
 We used all the artifacts as it was intended by the corresponding licenses. No personal information or offensive content is contained in the considered datasets.
-The original text of this paper was spell- and grammar-checked and slightly smoothed out using Grammarly.
-Appendix G Potential risks
+The original text of this paper was spell- and grammar-checked and slightly smoothed using Grammarly.
+Appendix H Potential risks
   1. 1.